### v1

# Optical-Cavity Manipulation Strategies of Conical Intersections Mediated Singlet Fission Systems

Abstract
We offer a theoretical perspective on simulation and engineering of polaritonic conical-intersection-driven singlet-fission (SF) materials. Using rubrene as an example and applying the numerically accurate Davydov-Ansatz methodology, we derive dynamic and spectroscopic responses of the system and demonstrate key mechanisms capable of SF manipulation, viz. cavity-induced enhancement/weakening/suppression of SF, population localization on the singlet state via engineering of the cavity-mode excitation, polaron/polariton decoupling, collective enhancement of SF. We outline unsolved problems and challenges in the field, and share our views on the development of the future lines of research. We emphasize the significance of careful modeling of cascades of polaritonic conical intersections in high excitation manifolds and envisage that collective geometric phase effects may remarkably affect the SF dynamics and yield. We argue that microscopic interpretation of the main regulatory mechanisms of the polaritonic conical-intersection-driven SF can substantially deepen our understanding of this process, thereby providing novel ideas and solutions for improving conversion efficiency in photovoltaics.
With the constant emergence of novel ideas, concepts, and breakthroughs, the boundaries between disciplines have become increasingly blurred, leading to the ascent of highly interdisciplinary fields. One such example is the intersection of molecular physics/chemistry, photonics, and materials science, which gave a birth to optical cavity molecular science and optical cavity materials science Ribeiroa ; Vidal21a ; Bhuyan . By harnessing the strong coupling between photonic modes of an optical cavity (or localized surface plasmon modes of nanoscale structures) and electronic/vibrational degrees of freedom (DOFs) of molecular systems, hybrid electron-vibration-photonic states can be formed Vidal22 ; Huo ; Uji23 ; Rubio23 and characterized by means of nonlinear ultrafast spectroscopy Fassioli ; Volodya23 .
Quantum electrodynamics (QED), the theoretical tool indispensable for modeling/understanding cavity-induced phenomena, was developed in the 1948-1949s Schwinger ; Feynman ; Tomonaga ; Dyson and recently heralded its 75th anniversary (see Orzel for a mini review). The polariton line was observed in the ZnO crystal in 1965 Thomas65 . The light-matter strong-coupling regime was demonstrated on phonons in the thin LiF film in 1975 Zhizhin75 and realized for a single Rydberg atom in the superconducting cavity in 1985 Mueller85 ; Klein87 . In the early 1990s, the strong coupling between excitons and photonic modes was experimentally demonstrated in semiconductor quantum dots Weisbuch . Since then, formation of polaritons owing to the strong coupling between optical-cavity/surface-plasmon modes and molecular excitons in various systems and under different scenarios have attracted significant attention, and a novel area of research known as polaritonic photophysics/photochemistry (PPP) has been established Ribeiroa ; Vidal21a ; Bhuyan ; Vidal22 ; Huo ; Uji23 ; Rubio23 ; Fassioli ; Volodya23 .
Figure 1: PPP-mediated SF. A collective bright singlet state coupled to the cavity mode forms the LP and UP states, while the collective singlet dark states are decoupled from the cavity mode. (a) Energy diagram illustrating the initial step of the SF process in the presence of collective strong coupling. The LP efficiently channels population towards the TT state. (b) A schematic illustration of the slanted rubrene microcavity Takahashi .
PPP deals with a large variety of processes, for example photoisomerization Fregoni ; Fregoni1 , photoinduced reactions in optical cavities or on metal substrates Yang ; Lather , electron and energy transfer processes Mandal ; Campos . Several intriguing phenomena have been discovered in the PPP studies. For instance, cavity-induced modification of the excitonic transport occurs mainly through the delocalized polariton modes rather than tunneling Schachenmayer . Cavities may enhance the exciton coherence length, as well as alter and even eliminate vibronic coupling in J-aggregates Spano15 ; Spano16 . Organic semiconductors showcase increasing conductivity in optical cavities Orgiu , and Raman scattering is enhanced owing to the presence of polaritons Mueller . Strong light-matter coupling has been experimentally demonstrated for chromosomes of photosynthetic bacteria Coles , photosynthetic complexes Tsargorodska , and even for living Chlorobaculum tepidum bacteria InVivo . These and other phenomena demonstrate that PPP holds potential for exciting practical applications, e. g., for engineering novel quantum materials CavityMaterials and development of polaritonic devices KC16 .
One of such applications, which holds a great promise for artificial light harvesting Artificial22 ; Artificial22a and can be substantially improved by the PPP machinery, is singlet fission (SF). It is a spin-allowed photophysical process discovered over fifty years ago SF1 , wherein a molecular material absorbs a visible-light photon generating a singlet exciton S which subsequently undergoes a spin-allowed transformation into two triplet excitons TT Smith1 ; Smith2 ; Rev1 ; Rev2 ; JCP_SF ; Rev3 . Nearly a decade ago, it was shown that dynamics of excitons can be altered drastically by embedding them in an optical cavity Schachenmayer . The idea of utilizing cavities to enhance SF was later proposed, drawing inspiration from the collective enhancement of SF efficiency observed in acene aggregates Mart . Then the pivotal role of the formally dark states for the amplification of SF in the strong-coupling regime in organic materials was demonstrated CavitySF .
Figure 1 sketches the essentials of the cavity-enhanced SF, showing the lower polaritonic (LP) state and the upper polaritonic (UP) state producing via the coupling of the SF dimer to the cavity. Considering the large density of states in the manifolds of dark and TT states, it is essential that the TT states have to be energetically lower than the dark states, in order to prevent population leakage and enhance the SF yield (panel a). Hence the LP state acts as a new channel towards the TT state to accelerate SF. Figure 1(b) depicts the optical microcavity, in which the rubrene amorphous film is sandwiched between two silver mirrors for producing polaritonic states Takahashi . By constructing the slanted rubrene microcavity, Takahashi et al. experimentally demonstrated the polaron decoupling effect Takahashi theoretically predicted by Spano Spano15 . They also found three critical parameters governing the cavity-mediated SF, the excitonic fraction in the LP state, the energy alignment between the LP and TT states, and the Frank-Condon factor for the excittaion Takahashi .
Table 1: Summary of bare SF material parameters and their optimal gain rates of TT yield $\tilde{\epsilon}_{\rm TT}/\epsilon_{\rm TT}$ in the cavity. where $\Delta\rm G$ is given by $\Delta\rm G=\omega_{\rm TT}-\omega_{\rm S_{1}}$, and $\epsilon_{\rm TT}$ ($\tilde{\epsilon}_{\rm TT}$) represents bared (polariton-assisted) TT yield, respectively.
$\rm Molecule$
$\Delta\rm G(meV)$
$\epsilon_{\rm TT}(\%)$
$\tilde{\epsilon}_{\rm TT}/\epsilon_{\rm TT}$
Ref.
$\rm Tetracene$
$150$
0.01
$\sim 16$
Mart
$\rm Pentacene$
$-110$
100
$\sim 1.4$
Mart
$\rm Hexacene$
$-630$
48
$\sim 3.3$
Mart
$\rm Rubrene$
$50$
17,40,>99.917,40,>99.917 , 40 , > 99.9
$\sim 1.3$
Takahashi ; Maslennikov ; Malin13
The parameters characterizing several popular bare and cavity-manipulated SF materials are listed in Table 1. As follows from this Table, incorporation of these materials in microcavities increases SF yield. Hence integration of QED-cavities into the design of photovoltaic devices offers a notable potential to enhance singlet-to-triplet conversion efficiency and circumvent the Shockley-Queisser limit SQ ; Markvart22 . However, the fulfilment of this program requires understanding of the microscopic mechanisms of the cavity-enhanced SF. In this context, it is essential that the majority of the SF materials (for example, pentacene Musser ; HGD20 , tetracene Herbert17 ; Xiao20 , and rubrene Miyata ) exhibit conical-intersection (CI) driven SF.
CIs are points of degeneracy in the potential energy surfaces (PESs) of isolated polyatomic molecules. Manifolds of these points of degeneracy, known as CI seams, undermine the Born-Oppenheimer approximation and open up effective and fast channels for the population and charge transfer ConicalIntersections . In PPP systems, the situation is complicated by the fact that CIs may be either intrinsic or cavity-induced Mukamel16 ; Kowalewski17 . In the strong coupling regime, an intrinsic molecular CI splits into a pair of polaritonic CIs Gu2 ; sun1 . Furthermore, strong coupling to the cavity mode may produce so-called light-induced CIs Cederbaum12 ; Cederbaum13 , which occur even in diatomic molecules possessing no intrinsic CIs Cederbaum08 ; Csehi . These light-induced CIs may be caused by the presence of additional (for example, rotational/orientational) DOFs. Hence the coupling to the cavity creates multidimensional PES landscapes comprising cascades of intrinsic and cavity-induced/modified CIs Sun6 ; Gu3 ; Gu2 ; sun1 ; Cho ; Gu ; Gu1 ; sun2 . On the one hand, this enormously complicates the situation and imposes high requirements on the numerical efficiency and accuracy of the theoretical methods used for the simulation of dynamic and spectroscopic responses of these systems. On the other hand, the cavity-induced/modified CIs create additional pathways and shortcuts for the population transfer, which can be used for engineering/optimizing the SF process Gu3 ; sun1 . Yet, the polaritonic CIs in the strong-coupling regime may act as double-edged sword. For example, Gu et al. uncovered that SF can be significantly suppressed, since polaritonic CIs are shifted away from the Franck-Condon region in pentacene dimers Gu2 . Hence enhancing the SF process in cavities remains a topic of contention.
In this Perspective, we summarize the status quo of the studies into the CI-driven cavity-mediated SF from the point of view of theoreticians. We begin with the overview of the fundamental processes shaping the SF mechanisms and outline spectroscopic signatures of these mechanisms. Then we put these findings in the general context of theoretical studies of PPP molecular systems and materials, outlining the open questions and computational challenges. For making the present work reader-friendly, we choose to concentrate on a single representative SF system, viz. rubrene dimers (RDs). All simulations of the present work were performed with the multiple Davydov Ansatz (mDA) method zhao2 ; zhao1 which is a well-established many-body wave-function-based variational integrator of quantum dynamics employing Gaussian coherent states (see the last section of this work for the discussion of the suitable numerical tools for the simulation of PPP dynamics and spectroscopy). If the mDA multiplicity is high enough (this is so for all numerical illustrations of the present work), the mDA can be regarded as the numerically accurate method which faithfully reproduces evolutions of quantum systems with many DOFs.
Polaritonic CI landscapes. In rubrene, the energy of the TT state, $\rm E_{TT}$ is slightly higher than the energy of the singlet state, $\rm E_{S_{1}}$. Hence SF – without thermal activation Miyata or without excitation to higher-lying electronic states $\rm E_{S_{n}}$ Malin12 – is rather inefficient. Furthermore, the usual mechanism of polaritonic enhancement of SF through the cavity-finetuned resonance of the LP and TT states, is ineffective in rubrene, too Ribeiroa ; Takahashi ; Mart . Hence rubrene is a suitable system for exploring novel cavity-induced mechanisms of the SF enhancement.
Figure 2: Cuts of the adiabatic PESs along the tuning mode ($Q^{(j)}_{c}=0$) for one (left column; panels a, b) and two (right column; panels c, d) RDs in the cavity for the first few polaritonic manifolds $N$. (a) No coupling to the cavity mode ($\Omega=0$), the bare CI is located in the Franck-Condon region at $Q^{(1)}_{t}=0.07,E=2.256\leavevmode\nobreak\ \rm eV$. (b, c, d) Strong coupling to the cavity mode ($\Omega=0.2\leavevmode\nobreak\ \rm eV$). The fraction of the TT states ($W_{\rm TT}$ of Eq. (3)) is indicated through the line color coding of PESs (a, c), while fraction of the photonic states ($W_{\rm ph}$ of Eq. (2)), is indicated through the line color coding of the PESs (b, d). The dash lines in panel (b) show positions of the polaritonic CIs.
Following Ref. sun1 , we consider a microcavity with one or two RDs placed into it. To focus on the net effect of the cavity-induced control mechanisms, the loss of photons in the cavity is neglected. This is justified, since the ultrahigh-Q PhCnB cavities ($\rm Q\sim 750,000$) have been created by using a five-hole taper design Deotare and the timescale of the studied processes (hundreds of femtoseconds) is quite short.
In this section, the $j$th RD participating in the CI-mediated SF is modeled as a system with three electronic states (the ground state ${\rm S}_{0}^{(j)}$, the singlet excited state ${\rm S}_{1}^{(j)}$, and the correlated triplet pair state ${\rm TT}^{(j)}$) and two vibrational modes (the tuning mode with a frequency $\Omega_{t}=0.186$ eV and coordinate $Q^{(j)}_{t}$, and the coupling mode with a frequency $\Omega_{c}=0.0154$ eV and coordinate $Q^{(j)}_{c}$iyata ; sun1 . The electronic energies of the singlet and triplet states are fixed at $E_{{\rm S}_{1}}^{(j)}=2.23$ eV and $E_{{\rm TT}_{1}}^{(j)}=2.28$ eV. The cavity is modelled via a single photonic mode of a frequency $\omega_{c}=2.256$ eV, which is nearly in the electronic resonance with the ${\rm S}_{1}^{(j)}$ and ${\rm TT}_{1}^{(j)}$ states. $\Omega$ determines the coupling of the RDs to the cavity mode. If not explicitly stated otherwise, we set $\Omega=0.2$ eV, which corresponds to the strong cavity-RD coupling regime. To increase space of the adjustable parameters, the rotating-wave approximation (RWA) in the description of the cavity-matter interaction was not applied.
All levels of the combined RD(s) + cavity system can be classified according to the expectation value of the number operator
$N=\langle\Psi_{n}({\bf Q})|\hat{N}_{\rm ex}|\Psi_{n}({\bf Q})\rangle$ (1)
where $|\Psi_{n}({\bf Q})\rangle$ is the adiabatic eigenfunction of the polaritonic Hamiltonian. $N$ specifies the excitation manifold of the polaritonic PESs. For example, $N=0$ corresponds to the collective ground state, $N=1$ yields the singly-excited manifold, etc. To characterize and visualize various states, it is elucidating to introduce several quantifiers. For instance,
$W_{\rm ph}=1-\sum_{\boldsymbol{k}}\sum_{n_{c}\neq 0}|\langle\boldsymbol{k},n_{c}|\Psi_{n}({\bf Q})\rangle|^{2}$ (2)
gives a fraction of the pure photonic excitation in a certain manifold $N$. The sum in Eq. (2) runs over the components $k_{j}={\rm S}^{(j)}_{1},{\rm TT}^{(j)}$ of vector $\boldsymbol{k}$. The number of these components matches the number of RDs in the cavity, and $n_{c}$ is the quantum number specifying the photonic-mode excitation. Similarly, the TT component of the polaritonic state can be defined as
$W_{\rm TT}=\sum_{\boldsymbol{k}\in{\rm TT}}\sum_{n_{c}}|\langle\boldsymbol{k},n_{c}|\Psi_{n}({\bf Q})\rangle|^{2},$ (3)
where the notation $\boldsymbol{k}\in{\rm TT}$ means that at least one of the components $k_{j}$ must contain ${\rm TT}^{(j)}$.
The fractions of the TT states ($W_{\rm TT}$ of Eq. (3)) and photonic states ($W_{\rm ph}$ of Eq. (2)) give a convenient visualization of the polaritonic PESs, which is implemented through the line color coding in Figure 2. Panels (a, b) represent a single RD in the cavity, and show the cut of the polaritonic PESs along the symmetric tuning mode $Q_{t}$ at $Q_{c}=0$. Panel (a) corresponds to a cavity-free RD. Hence, ${\rm S}_{1}^{(1)}$ and ${\rm TT}^{(j)}$ are not dressed with the cavity mode, and only the intrinsic RD CI at $Q_{t}=0.07,E=2.256\leavevmode\nobreak\ \rm eV$ exists in this case. It follows that this CI located in the Franck-Condon region provides the only SF pathway. In the presence of cavity, the PESs change substantially, and cascades of polaritonic CIs appear in Figure 2(b). As the coupling to the cavity increases, the two polaritonic CIs move away from the Franck-Condon point, the distance between the pair of polaritonic CIs increases with $N$, and positions of these CIs lie along the V-shaped dash lines in Figure 2(b). It is essential that the fraction of the hybrid matter-photon PESs increases with $N$ near the Franck-Condon point. This facilitates the cavity-assisted SF.
The cuts of PESs for two RDs strongly coupled to the cavity mode are shown in Figures 2(c) (highlighting the TT character of the PESs) and (d) (highlighting the photonic character of the PESs). First of all the number of the coupled PESs apparently increases with the number of RDs in the cavity. If for a single RD we have 3 PESs for $N\geq 1$, for two RDs there are 5 PESs for $N=1$, and 9 PESs for $N\geq 2$. The PES with the CIs at $Q^{(1)}_{t}=Q^{(2)}_{t}=0.07$ eV, which corresponds to both RDs in the triplet state, exists for all $N$. With the exception of this PES, the cavity-RD coupling repels polaritonic CIs from the central region, but the effect is not so pronounced as in the case of a single RD.
SF Dynamics Mediated by Polaritonic CIs. Let us explore how the polaritonic PESs coupled through cascades of the cavity-induced CIs affect the SF dynamics. For this to achieve, it is worthwhile to study what is the impact of the initial excitation of the photonic mode on the ensuing singlet-to-triplet population transfer. In our simulations, the cavity mode is assumed to be initially, at $t=0$, in the coherent state
$|\mu_{1}(t)\rangle=\exp(-|\mu_{1}|^{2}/2)\sum_{n_{c}=0}^{\infty}\frac{\mu_{1}^{n_{c}}}{\sqrt{n_{c}!}}|n_{c}\rangle$ (4)
where $\mu_{1}$ is the displacement parameter. For a specific $n_{c}$, the main contribution to the coherent state gives the term with $\mu_{1}(0)=\sqrt{n_{c}}$. Hence we will systematically study how the initial preparation of the RD system in the manifold $N$, where the population is initially placed in the singlet state and the cavity mode is in the coherent state with $\mu_{1}(0)=\sqrt{N}$ affects SF in the regime of strong RD-cavity coupling. This is illustrated by Figure 3 for one (a, c) and two (b, d) RDs in the cavity.
Figure 3: Population dynamics of the triplet state (upper panels) and average photon mode populations $P_{cav}(t)$ (lower panels) for several initial photon-mode pumping numbers $|\mu_{1}(0)|^{2}=N$. Left column (panels a, c): one RD. Right column (panels b, d): two RDs.
Let us begin with the considerations of the upper panels, which show the triplet populations for different $N$. For a single RD, the triplet population – on the average – increases with $N$ up to $N=6$. For example, the population of $\approx 0.28$ is observed at 300 fs for $N=6$, which is approximately twice larger than the population obtained in the same model without a cavity. However, the SF efficiency starts to decrease at $N=7$ and is fully suppressed for $N\geq 8$. This effect can rationalized from the different points of view. On the one hand, it can be viewed as excitation localization on the singlet state ${\rm S}_{1}^{(1)}$ and on the photonic mode, which renders the ${\rm TT}^{(1)}$ effectively decoupled and the SF switched off (see Refs. Localization19 ; Localization20 for engineering of long-lived localized states in the cavity-QED systems). A qualitatively similar localization is realized, for example, in the spin-boson model in the regime of the strong coupling of the spin system to the bosonic bath Wu , even at different system preparations LP23 . Alternatively, the effect can be envisaged as an effective annihilation of the nonadiabatic coupling between the singlet and triplet states Gelin09 . It operates through the combined effect of the cavity field (through Rabi cycling) the system itself (through the intrinsic ${\rm S}_{1}^{(1)}-{\rm TT}^{(1)}$ CI), since the photonic mode does not directly couple the singlet and triplet states. Note also that the population evolution of the triplet states starts not from $t=0$, but from a characteristic time $\tau_{N}$ which increases with $N$. This indicates that the wavepacket needs $\approx\tau_{N}$ to reach the polaritonic CIs.
[…]
Let us now inspect evolutions $P_{cav}(t)$ of the cavity-mode populations vs. the initial photon pumping numbers $N$ ($P_{cav}(0)=N$) which deliver a complementary information. $P_{cav}(t)$ for a single RD is depicted in Figure 3(c). For $N$ from 1 to 6, the populations exhibit similar behavior: During the first 50 fs, they show several oscillations with a rough period of 22 fs which is related to the tuning mode. Then they reach a quasi-steady-state $\langle P_{cav}\rangle$ value at longer times. It is worth mentioning that $\langle P_{cav}\rangle$ are slightly larger than the initial populations $P_{cav}(0)=N$. It reveals the population redistribution due to the coupling of the cavity mode and the ${\rm S}_{1}^{(1)}$ state. $P_{cav}(t)$ for $N=7$ shows high-amplitude oscillations during the first $\sim 40$ fs, which matches the time $\tau_{7}$ during which the corresponding triplet population in Figure 3(a) is almost zero. For $N=8$, the high-amplitude oscillations persist over the entire evolution of $P_{cav}(t)$, which is a signature of the localization of the population on the photonic mode and the singlet state, which results in the SF stopping. Interestingly, the period of these fast oscillations reveals the effective Rabi frequency $\Omega_{N}\approx\Omega\sqrt{N}$ ($2\pi/\Omega_{8}\approx 7.3$ fs), which is defined under the assumption that the effective coupling to the cavity scales $\sim\sqrt{N}$.
Figure 3(d) is a counterpart of Figure 3(c), but for two RDs coupled to the cavity. For $N$ from 1 to 6, $P_{cav}(t)$ reach their first local minima at $t\approx 17$ fs (dashed line in panel d), which indicates that a fraction of the photonic-mode population is transferred to the doubly excited electronic states of the RDs. Then, $P_{cav}(t)$ exhibit oscillations with a period of $\sim 26$ fs, which cannot be directly related to the model parameters. In general, excitation to $1\leq N\leq 6$ reflects population transfer between photonic mode and excited states of the RDs. For $7\leq N\leq 9$, $P_{cav}(t)$ shows fast Rabi oscillations at the initial stage, which correlate with the absence of the triplet excitation at the same time interval (cf. Figure 3(b)). For $N=10$, fast Rabi oscillations spread over the entire $P_{cav}(t)$ evolution, which is the signature of the SF switching-off in Figure 3(b). In the long-time limit, the $P_{cav}(t)$ reach quasi-steady states $\langle P_{cav}\rangle$ the values of which are lower than the initial values $P_{cav}(0)$. This indicates a net population transfer from the photonic mode to the doubly-excited electronic states of RDs.
Figure 4: The level structure for two rubrene dimers in the cavity. The superscripts $(1)$ and $(2)$ number the dimers.
The above analyses demonstrate that the SF efficiency is primarily determined by two factors. The first factor is the position of the polaritonic CIs (see Figure 2). If the CIs are located in the Franck-Condon region, the subsequent population transfer is facilitated and the higher SF efficiency is achieved. The second factor is the initial wavepacket preparation and the state of the cavity which is controlled by the photonic-mode pumping number $|\mu_{1}(0)|^{2}=N$. This later parameter plays the crucial role, increasing the SF efficiency for small $N$ and fully suppressing SF once $N$ reaches a certain critical value.
Spectroscopic Signatures of Polaritonic CIs. As demonstrated in the previous sections, strong coupling of RDs to the cavity mode drastically alters the SF efficiency owing to the cohesive action of the cavity-induced and intrinsic CIs. This processes are imprinted into nonlinear femtosecond spectroscopic responses of these systems. To find spectroscopic signatures of the polaritonic SF occurring in multidimensional PES landscapes, the signals must be accurately simulated and adequately interpreted. For this to achieve, it is essential to realistically model RDs. Indeed, all third-order spectroscopic responses consist of three contributions, the ground-state bleach (GSB), the stimulated emission (SE), and the excited-state absorption (ESA) Darius04 ; Leonas . The GSB and SE reflect the photoinduced processes in the manifolds with $N=0$ (the ground state) and $N=1$ (singly-excited states), while ESA involves also the doubly-excited states, $N=2$. Hence modeling individual RDs as systems containing only singly-excited states ${\rm S}_{1}^{(j)}$ and ${\rm TT}^{(j)}$ is insufficient: we have to include higher-excited states ${\rm S}_{n}^{(j)}$ and ${\rm TT}_{n}^{(j)}$ sun2 . Following Miyata ; sun2 , we set $E_{{\rm S}_{n}}^{(j)}=4.33$ eV and $E_{{\rm TT}_{n}}^{(j)}=4.68$ eV in our simulations. For two RDs in the cavity, we thus obtain a single ground state, 5 singly-excited states and 15 doubly-excited states which are specified in Fig. 4. The states, in turn, can be subdivided into the matter states ($n_{c}=0$, left part of Fig. 4) and hybrid states ($n_{c}>0$, right part of Fig. 4).
Here we consider electronic two-dimensional (2D) spectra, because electronic 2D spectroscopy (2DES) is one of the most powerful and information-rich third-order spectroscopic techniques (see Scholes22 and references therein). Qualitatively, 2DES spectra $S(\omega_{\tau},T_{w},\omega_{t})$ can be envisaged as a $T_{w}$-dependent 2D $(\omega_{\tau},\omega_{t})$ maps, which reveal correlations between the states of the system excited at a frequency $\omega_{\tau}$ at time zero and probed at a frequency $\omega_{t}$ at time $T_{w}$. 2DES spectra $S(\omega_{\tau},T_{w},\omega_{t})$ simulated for the two RD system strongly coupled to the cavity mode are presented in Figure 5 for different population times $T_{w}$. From top to bottom, shown are the SE (row (a)), GSB (row (b)), and ESA (row (c)) contributions, as well as the total signal (row (d)).
The SE contribution (row (a)) possesses two pronounced diagonal peaks (DPs) and two cross-peaks (CPs) (similar situation takes place in the purely polaritonic dimer system Bondarev13 ). The DP1 ($\omega_{\tau}=2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$) and DP2 ($\omega_{\tau}=2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) reveal the middle polariton (MP) and the UP bands, respectively. The low-intensity LP band cannot be reliably identified in the SE 2DES spectrum, because 2DES peak intensities are proportional to the product of four relevant transition dipole moments, in contrast with absorption spectra, which are proportional to the product of two transition dipole moments (see the discussion in Ref. Jindra15 ). The CP1 ($\omega_{\tau}=2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$) and CP2 ($\omega_{\tau}=2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) quantify the coupling and polaritonic coherence between the states MP and UP. A weak CP ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) at $T_{w}=16$ fs reveals the phonon sideband. The shape and intensity of the DPs and CPs depend significantly on the waiting time $T_{w}$. The DPs at $T_{w}=0$ are predominately elongated perpendicular to the main diagonal, but their elongation turns into parallel at longer population times. A certain similarity of the DPs at $T_{w}=0$ and 32 fs shows a significant contribution of the polaritonic beating with a period of $\sim 32\leavevmode\nobreak\ {\rm fs}$ revealing the energy difference of $0.132\,{\rm eV}$ between the states MP and UP. The two negative CPs are elongated parallel to the main diagonal. Similar to the DPs, the CPs are also significantly modulated by the MP-UP energy-gap beating. The overall decreases of the intensity of the SE contribution with population time mirrors the CI-mediated population transfer from the bright $S_{1}^{(j)}$ states to the dark $TT^{(j)}_{1}$ states and reflects thereby the SF dynamics.
The GSB contribution is shown in panels (b) of Figure 5. It reveals the same pairs of the polaritonic DPs and CPs and, additionally, a pair of weaker polariton-phonon CPs, ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) and ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$). In comparison with the SE peaks, their GSB counterparts look rather static. Nevertheless, the GSB contribution reveals the vibrational wave-packet motion in the electronic ground state, which is clearly manifested in the peak evolutions (see Ref. sun2 ).
The ESA contributions are depicted in panels (c) of Figure 5. The SE and ESA contributions to 2DES spectra reveal photophysical processes in the manifold $N=1$ skw20a . However, the two contributions employ different reporter states for delivering the same information: the SE employs the electronic ground state $N=0$, while the ESA employs the doubly-excited manifold $N=2$. Hence shapes and evolutions of predominantly positive SE contributions and predominantly negative ESA contributions are quite similar. Nevertheless, positions of the DPs and CPs in the SE and ESA spectra are slightly shifted with respect to each other. At $T_{w}=0\leavevmode\nobreak\ {\rm fs}$, for example, the (nearly) diagonal ESA peaks are located at ($\omega_{\tau}\approx 2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.317\leavevmode\nobreak\ {\rm eV}$) and ($\omega_{\tau}\approx 2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.416\,{\rm eV}$). All ESA peaks are elongated parallel to the $\omega_{t}$ axis, due to a high density of states in the manifold $N=2$. The peak ($\omega_{\tau}\approx 2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.317\leavevmode\nobreak\ {\rm eV}$), which revels transitions from the MP state to the states of the doubly-excited manifold, has the highest intensity. Note also that the relative separation between the DPs and CPs along the $\omega_{t}$ axis changes with $T_{w}$: cf. the spectra at 0 fs with those at 16 fs. This modulation is caused by the superposition of polaritonic and purely vibrational tuning-mode oscillations. This reflects a higher complexity of the ESA contribution, which involves polaritonic states with $N=1$ and 2.
Finally, the total 2DES signal $S(\omega_{\tau},T_{w},\omega_{t})$ is shown in Figure 5(d). Since positions of the DPs and CPs in the SE, GSB and ESA contributions nearly coincide, the total 2DES spectrum exhibits positive, negative and dispersed peak shapes, the amplitudes of which are determined by subtle cancellation effects in the superposition of positive SE and GSB contributions and negative ESA contribution. The overall structure of the total 2DES spectra is dominated by the quartet of polaritonic DPs and CPs. In addition, a number of polariton-vibrational peaks around $\omega_{\tau}\,\mathrm{or}\,\omega_{t}=$ 2.6 eV is seen at different $T_{w}.$ It is remarkable that these hybrid polariton-vibrational peaks are more pronounced in the total signal than in the SE, GSB, and ESA contributions – owing to the aforementioned cancellation effects. The DPs are positive while the CPs exhibit unusual asymmetry: the upper CP is always positive, but the spectral structures around the lower CP can be negative (see the signals for $T_{w}=0$ and 16 fs).
Figure 5: Real part of the SE (row (a)), GSB (row (b)), ESA (row (c)) contributions and the total 2DES spectrum (row (d)) of the two RD system for several population times $T_{w}$ indicated in the panels. The spectra are computed in the impulsive limit.
Brief summary. We have applied a numerically accurate mDA methodology to explore dynamic and spectroscopic responses of the CI-driven SF RD system strongly coupled to the cavity mode. Several microscopic SF mechanisms have been uncovered, viz. proximity of the CI-driven polaritonic wave packet to the Franck-Condon region, cavity-induced enchantment/weakening/suppression of SF, polaron/polariton decoupling, localization of the singlet-state population via engineering of the cavity-mode excitation, collective enhancement of SF.
Simulation and inspection of nonlinear femtosecond spectroscopic signals of the polaritonic CI-SF systems provides accurate information on the singlet-to-triplet transfer in real time. In particular, 2DES spectra of the RD system are dominated by the quartet of DPs and CPs that reveal the MP and UP bands, while the hybrid polariton-vibrational CPs contain information on the UP sideband. This clear separation of the polaritonic and vibrational peaks is a noticeable feature of the obtained spectra. In addition the simulated 2DES spectra exhibit multiple peak-cancellation effects, and their population-time evolution is accompanied by polariton-vibrational beatings.
We conclude that microcavities can open up new avenues for improving efficiency and stability of the CI-mediated SF, and ultrafast nonlinear spectroscopic signals – if interpreted with adequate theoretical support – can provide the real-time guidance on the SF reaction pathways. Yet, the polaritonic SF studies have not reached its maturity, and a lot of problems await efficient and practical solutions. The specifics are detailed below.
Open questions, challenges, and perspectives.
System Hamiltonian. In excitonic systems, doubly-excited states are usually understood as the states in which two chromophores are simultaneously excited. This interpretation, however, is not always sufficient for SF aggregates, because polyatomic molecules comprising these aggregates possess dense manifolds of higher-lying excited electronic states and the minimal system supporting the SF reaction is not a single chromophore, but a dimer. Usually, higher-lying electronic states do not contribute to the SF per se (see the discussion below though). Yet, optical transitions between the lower-lying and higher-lying states affect dramatically ESA contributions to spectroscopic signals. Hence higher-lying states of individual molecules comprising SF dimers, aggregates, and materials have to be taken into account in realistic simulations of spectroscopic responses.
It is useful to characterize all relevant electronic states and parameterize the SF Hamiltonians by using ab initio methods of quantum chemistry MT18 ; MT19 ; YiZhao21 ; Haibo21 ; Vendrell23 . In addition, ab initio methods can be applied to the entire system of molecules/aggregate+cavity. For example, quantum mechanics/molecular mechanics (QM/MM) simulations of Refs. Gerrit22 ; Gerrit22a treat up to 512 polyatomic chromophores in the cavity. Simulations based on the ab initio QED Rubio23 ; Koch21 , multiple spawning Martinez24 , and cavity Born-Oppenheimer Hartree-Fock Ansatz Kowalewski23 can also provide valuable insight into SF reactions in specific systems.
Quantum dynamic and spectroscopy. Up to now, accurate quantum simulations of up to two CI-carrying molecules/dimers in the cavity have been performed Gu1 ; sun1 ; sun2 . It is of great interest to push the simulations towards higher number of molecules/dimers $N_{M}$ and explore how CIs affect cooperativity, e.g., to check the conventional $\sim\sqrt{N_{M}}$ scaling of the effective system-cavity coupling Huo22 . For this to achieve, few-molecule models with symmetry-optimized Hamiltonians JYZ23a as well as multi-configuration time-dependent Hartree (MCTDH) MT18 ; MT19 ; Vendrell19 ; Vendrell20 , tensor-network Vidal18 ; Keeling22 , and mDA methods look promising. Recently emerging emulations of CIs on quantum simulators CIsimulations23a ; CIsimulations23b ; Schoelkopf23 may also be extended in the future to cavity-CI systems.
Disorder is of paramount importance for realistic modeling of CI-driven cavity-mediated SF systems, taking into account that defects/disorder can induce CIs in semiconducting molecular materials Chuntonov22 and affect significantly polaritonic transport Schachenmayer22 ; HTC . Temperature is essential, e.g., for activating CI-mediated SI in rubrene Miyata and for catalysing reactions in polaritonic systems Campos . The mDA protocol based on the thermo field dynamics (TFD) BG21 was recently applied to explore dynamics of the pristine RD skw ; Sun20 and the Holstein-Tavis-Cummings model Hou23 at finite temperature.
Apart from mDA, MCTDH and variants of tensor-network methods are especially favorable for performing fully-quantum simulations of realistic polaritonic CI-driven SF systems, because they permit the efficient evaluation of nonlinear spectroscopic signals Santoro23 ; GB_JCTC and – by using the TFD methodology BG21 and the recipe of Ref. GB21 – finite temperature and static disorder can also be taken care of in the (relatively) computationally inexpensive manner.
An extra challenge is to identify spectroscopic signatures of polaritonic effects in nonlinear spectroscopic signals of polaritonic CI-driven SF systems. Hence accurate theoretical simulations and interpretations of the existing Takahashi ; Watanabe20 ; Chergui20 and forthcoming experiments are demanding but rewarding.
Microscopic mechanisms behind polaritonic CI-driven SF. There exist several methods that can potentially enhance the SF yield. The conventional practices are based on the adjustment of the cavity mode for achieving resonance and strong coupling of the LP or UP states with the triplet state or on the construction of the cavity-assisted SF pathways. On the other hand, the geometric-phase effects may also be harnessed for engineering SF in CI-driven polaritonic systems. The controlling mechanisms may be implemented through collective topological polaritonic phases JYZ23 that are akin to those in topological metamaterials Alu23 . It is also worthwhile to explore interfacing of light-induced CIs in cavities Vibok22a ; Vibok22b ; Huo21 with intrinsic CI systems, as their combination may facilitate fine-tuning of the multiple CI landscapes VendrellVibok .
Apart from the aforementioned cavity-related mechanisms, there exist other handles for tweaking SF, that are based on the optimization of the bias voltage Xiong18 , temperature Miyata ; Xiong15 ; Wasielewski21 , carrier frequency BY22 and intensity Asahi16 ; Bardeen18a of the excitation pulse, or even static disorder Chavez ; Disorder24 .
Yet, it is now always clear whether the coupling to the cavity (perhaps in combination with other methods mentioned above) enhances or suppresses the SF yield. It is promising, therefore, to interface the existing protocols for the dynamic simulations of polaritonic CI-driven SF systems with quantum control routines Control1 ; Control2 , in order to find a robust way to comprehensive optimization of SF in cavities.
I TOC Graphic
Acknowledgments
The authors gratefully acknowledge the support of the Singapore Ministry of Education Academic Research Fund (Grant No. RG87/20). K. Sun would also like to thank the Natural Science Foundation of Zhejiang Province (Grant No LY18A040005) for partial support. M. F. G. acknowledges support from the National Natural Science Foundation of China (Grant No. 22373028).