### v2

# Optical-cavity manipulation strategies of singlet fission systems mediated by conical intersections: insights from fully quantum simulations

Abstract
We offer a theoretical perspective on simulation and engineering of polaritonic conical-intersection-driven singlet-fission (SF) materials. We begin by examining fundamental models, including Tavis-Cummings and Holstein-Tavis-Cummings Hamiltonians, exploring how disorder, non-Hermitian effects, and finite temperature conditions impact their dynamics, setting the stage for studying conical intersections and their crucial role in SF. Using rubrene as an example and applying the numerically accurate Davydov-Ansatz methodology, we derive dynamic and spectroscopic responses of the system and demonstrate key mechanisms capable of SF manipulation, viz. cavity-induced enhancement/weakening/suppression of SF, population localization on the singlet state via engineering of the cavity-mode excitation, polaron/polariton decoupling, collective enhancement of SF. We outline unsolved problems and challenges in the field, and share our views on the development of the future lines of research. We emphasize the significance of careful modeling of cascades of polaritonic conical intersections in high excitation manifolds and envisage that collective geometric phase effects may remarkably affect the SF dynamics and yield. We argue that microscopic interpretation of the main regulatory mechanisms of polaritonic conical-intersection-driven SF can substantially deepen our understanding of this process, thereby providing novel ideas and solutions for improving conversion efficiency in photovoltaics.
I INTRODUCTION
Cavity molecular science is a rapidly developing interdisciplinary field that has gained significant attention in recent yearsGalego ; Ribeiro2018 ; Mandal . It explores the interaction between molecules and confined electromagnetic fields, such as those within optical or plasmonic cavities. By leveraging the strong coupling between light and matter and the highly hybridized electron-photon-phonon states thus formed, this field aims to manipulate chemical reactions, energy transfer processes, and molecular dynamics in novel manners. Applications have been found in diverse scenarios such as photo isomerization, singlet fission (SF), catalysis enhancement, selective bond manipulation, and intracavity electron transfer processes Orgiu ; Shalabney1 ; Riedinger ; Bienfait ; Vendrell . A typical example is the mixed light-matter quasi particles formed when the interaction between excitons and photons is strong enough, known as polaritons. In 1997, Agranovich et al. proposed a theory of strong light-matter coupling in organic molecules Agranovich , and in 1998, Lidzey et al. conducted the first experimental verification of this theory LIDZEY1 ; LIDZEY2 . In 2011, Schwartz et al. Schwartz introduced a photochemical coupling mechanism for optical switches under strong light, marking the beginning of advancements in polariton optics. Recent research has revealed that the strong coupling between molecular vibrations and a cavity mode can modify the rate constant of ground-state reactions, even in the absence of external driving. A notable characteristic observed in these experiments is the alteration of the vibrational spectrum, attributed to the formation of molecular vibrational polaritons – a hybrid state arising from the interaction between molecular cavity modes and vibrational modes. The existence of vibrational polaritons and changes in chemical reactivity can be characterized by strong resonance Kowalewski23 . Thus, controlling light-matter coupling can be leveraged for polariton transport engineering and the optimization of various photo-induced reactions Vendrell ; JYZ23a .
Figure 1: PPP-mediated SF. A collective bright singlet state coupled to the cavity mode forms the LP and UP states, while the collective singlet dark states are decoupled from the cavity mode. (a) Energy diagram illustrating the initial step of the SF process in the presence of collective strong coupling. The LP efficiently channels population towards the TT state. (b) A schematic illustration of the slanted rubrene microcavity Takahashi .
Six decades ago, Jaynes and Cummings drew inspiration from studies on quantum coherence in nuclear magnetic resonance and the solution to the Schrödinger equation for two-level systems in a rotating magnetic field to explain deviations between Fermi’s Golden Rule and perturbation theory in microwave resonators. This aligned with Rabi’s findings in 1937. As a result, the renowned Jaynes-Cummings (JC) model JAYNES was developed, describing the interaction between a single two-level atom and a single-mode cavity field under the rotating-wave approximation (RWA). Despite its simplicity, the JC model offers an elegant framework for understanding light-matter interactions. In recent years, various experimental setups have been devised to implement such a system BIRNBAUM ; MAUNZ ; WALLQUIST . Further research has revealed that the JC model exhibits significant complexity in terms of nonlinearity, leading to the development of various extended JC models. Among them, the most noteworthy are the Tavis Cummings (TC) model TAVIS and its variant without RWA, namely, the Dicke model DICKE . One of the most representative features of these Hamiltonians is a second-order quantum phase transition (QPT) from the normal phase to the super-radiant phase in the thermodynamic limit BAUMANN ; KLINDER ; BADEN , a collective phenomenon characterized by condensation of a macroscopic number of photons or atoms.
The interaction between the system and vibrational degrees of freedom (DOFs) can influence molecular behavior within the optical cavity, necessitating the incorporation of phonon modes into the TC model to create a composite framework. This composite model, known as the Holstein-Tavis-Cummings (HTC) model, is widely utilized to study various aspects of polaritonic transport. By tuning the interaction between atoms and cavity fields, steady-state exciton transport can be achieved, and factors influencing atomic transitions, as well as photon emission and absorption processes, can be examined. For the more complex HTC model, approximate analytical and numerical methods are typically employed to investigate its dynamic properties LIU ; Takahashi ; WU ; Spano15 ; Herrera2016 ; HERRERA2 ; HERRERA3 ; Cwik ; HERRERA4 .
Strong light-matter coupling has been experimentally demonstrated for chromosomes of photosynthetic bacteria Coles , photosynthetic complexes Tsargorodska , and even for living Chlorobaculum tepidum bacteria InVivo . These and other phenomena demonstrate that polaritonic photophysics and photochemistry (PPP) holds potential for exciting practical applications, e.g., for engineering novel quantum materials CavityMaterials and development of polaritonic devices KC16 . One of such applications, which holds a great promise for artificial light harvesting Artificial22 ; Artificial22a and can be substantially improved by the PPP machinery, is SF, a spin-allowed photophysical process discovered over fifty years ago in which a single photoexcited singlet exciton is converted into two triplet excitons (TT) SF1 ; Smith2010 ; Smith2 ; Rev1 ; Rev2 ; JCP_SF ; Rev3 ; Musser ; HGD20 ; Miyata . By effectively doubling the number of generated excitons from one photon, SF offers a pathway to exceed the Shockley-Queisser limit and thus enhance power conversion efficiency in solar cells Rao2017 ; Smith2010 ; Congreve2013 . As an all-organic and cost-effective approach to broaden spectral utilization, SF has stimulated extensive research aimed at understanding its underlying principles, optimizing its yields, and integrating it into diverse molecular architectures and solid-state platforms Berkelbach2013 ; Margulies2016 . Nonetheless, controlling the rate and coherence of SF remains challenging, partly due to the intricate interplay of electronic interactions, vibronic couplings, and complex many-body dynamics Monahan2015 ; Margulies2016 .
To transcend these molecular-level constraints, new strategies beyond chemical modification are being explored. One particularly promising route is to incorporate SF-active materials into optical cavities, drawing on concepts from polaritonic chemistry and cavity quantum electrodynamics Mart ; Schachenmayer ; Ebbesen2016 ; Ribeiro2018 ; Flick2017 ; Hertzog2019 . Embedding SF systems inside cavities forms polaritons hybrid light-matter states that reshape exciton energy landscapes and relaxation channels Munkhbat2018 ; Herrera2016 ; Galego2017 . Figure 1 sketches the essentials of the cavity-enhanced SF, showing the lower polaritonic (LP) state and the upper polaritonic (UP) state producing via the coupling of the SF dimer to the cavity. Considering the large density of states in the manifolds of dark and TT states, it is essential that the TT states have to be energetically lower than the dark states, in order to prevent population leakage and enhance the SF yield (panel a). Hence the LP state acts as a new channel towards the TT state to accelerate SF. Figure 1(b) depicts the optical microcavity, in which the rubrene amorphous film is sandwiched between two silver mirrors for producing polaritonic states Takahashi . By constructing the slanted rubrene microcavity, Takahashi et al. experimentally demonstrated the polaron decoupling effect Takahashi theoretically predicted by Spano Spano15 . They also found three critical parameters governing the cavity-mediated SF, the excitonic fraction in the LP state, the energy alignment between the LP and TT states, and the Frank-Condon factor for the excitation Takahashi . Furthermore, through tuning cavity modes, detuning frequencies, and manipulating Rabi splittings, it becomes possible to influence electron-phonon couplings, reduce nonradiative losses, and reorient pathways critical to SF efficiency Vendrell3 ; Tempelaar2017 ; Simpkins2021 ; CavitySF .
Moreover, conical intersections (CIs) topological intersections of excited-state potential energy surfaces b21 ; b22 ; b23 ; DS play a crucial role in regulating SF by determining how excitons evolve and split Berkelbach2013 ; Monahan2015 ; Musser ; Miyata . Under strong light-matter coupling, CIs can be effectively modified or navigated to favor SF channels Gu2 ; sun1 . Such cavity-mediated modifications complement molecular tailoring: rather than relying solely on molecular packing or orbital engineering, one can now exploit photonic environments to dynamically adjust exciton dynamics, stabilize intermediate states, and potentially boost SF yields Hertzog2019 ; Ribeiro2018 .
This Perspective focuses on how fully quantum simulations of dynamic observables and nonlinear spectroscopic signals elucidate the role of CIs and optical cavities in governing SF processes. By capturing the interplay of electrons, nuclei, and photons within a single quantum framework, these simulations provide new insights into cavity-mediated SF manipulation. They point toward cavity architectures that guide SF along desired pathways, offer greater resilience against disorder and thermal effects, and open unprecedented avenues for pushing SF efficiency beyond what is achievable in standard conditions Vendrell3 . Ultimately, this integration of SF systems with cavities promises to inform both fundamental research and technological innovations, facilitating devices with improved energy conversion, increased stability, and enhanced coherence lifetimes.
The remainder of this Perspective is structured as follows. Sec. II introduces the fundamental models of light-matter coupling, namely the TC and the HTC Hamiltonian. In Sec. II A, the dynamics and absorption spectra of the disordered TC and HTC models are analyzed. Sec. II B examines the impact of photon loss on HTC model dynamics, while Sec. II C explores the effects of temperature on HTC model dynamics and absorption spectra using the thermo-field dynamics (TFD) approach. Sec. III offers a theoretical perspective on simulations, focusing on CI landscapes (Sec. III A), SF dynamics (Sec. III B), and two-dimensional spectra (Sec. III C), providing new avenues for engineering polarized CI-driven SF materials. Sec. IV highlights unresolved issues, challenges, and future directions in polaritonic system simulations. Finally, the conclusions and discussions are drawn in Sec. V.
II Dynamics of light-matter coupling models
To lay the groundwork for more realistic Hamiltonians that capture the essential physics of cavity-assisted SF processes, we start with basic models of light-matter interactions, namely the TC and the HTC Hamiltonians. By examining these fundamental models, exploring how disorder, non-Hermitian effects, and finite temperature conditions impact their dynamics, we set the stage for studying more complex models focusing on CIs and their role in SF. The Hamiltonian of the TC model can be written as
$\hat{H}_{\mathrm{\rm TC}}=\omega_{c}\hat{a}^{\dagger}\hat{a}+\sum_{n=1}^{N}\left[\omega_{n}\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}+g_{n}\left(\hat{a}^{\dagger}\hat{\sigma}_{n}^{-}+\hat{a}\hat{\sigma}_{n}^{+}\right)\right],$ (1)
where $\omega_{c}$ is the frequency of the cavity mode, $\omega_{n}$ is the leap frequency of the $n$th qubit, and $g_{n}\equiv\omega_{\mathrm{R},n}/\sqrt{N}$ is the coefficient of qubit-cavity coupling in the case of a single quantum bit $(n=N=1)$. It is simplified to the Rabi frequency $\omega_{\mathrm{R}}$ $\hat{a}^{\dagger}(\hat{a})$ denotes the photon generation (annihilation) operator, $\hat{\sigma}_{n}^{+}=\left|1_{n}\right\rangle\left\langle 0_{n}\right|\left(\hat{\sigma}_{n}^{-}=\left|0_{n}\right\rangle\left\langle 1_{n}\right|\right)$ is the raising (lowering) Pauli operator on the two-energy level atoms. The Dicke model without the RWA is given by
$\hat{H}_{\mathrm{\rm Dicke}}=\omega_{c}\hat{a}^{\dagger}\hat{a}+\sum_{n=1}^{N}\left[\omega_{n}\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}+g_{n}\left(\hat{a}^{\dagger}+\hat{a}\right)\left(\hat{\sigma}_{n}^{-}+\hat{\sigma}_{n}^{+}\right)\right].$ (2)
The TC model and the Dicke model are very common tools for studying the strong coupling between atomic systems and cavity modes.
Vibrational DOFs play a vital role in accurately describing organic molecular cavity QED systems as the coupling between electronic states and lattice vibrations (phonons) is essential for understanding decoherence, emission properties, and the overall behavior of these systems. After adding a bath of phonons to $\hat{H}_{\mathrm{TC}}$, we arrive at the Hamiltonian for the HTC model
$\hat{H}_{\mathrm{HTC}}=$ $\hat{H}_{\mathrm{TC}}+\sum_{k}\omega_{k}\hat{b}_{k}^{\dagger}\hat{b}_{k}$
$-\frac{\lambda}{\sqrt{N}}\sum_{k}\sum_{n=1}^{N}\omega_{k}\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}\left(e^{-ikn}\hat{b}_{k}^{\dagger}+e^{ikn}\hat{b}_{k}\right),$ (3)
where $\omega_{k}$ and $\hat{b}_{k}^{\dagger}$ are the frequency and the production (annihilation) operator of the $k$th mode of the phonon, respectively. $\lambda$ is the diagonal coupling strength of the qubit-phonon.
II.1 Dynamics of disordered TC and HTC models
The TC model is analytically solvable since the total number of excitations $\hat{N}_{\rm ex}=\hat{a}^{\dagger}\hat{a}+\sum_{n=1}^{N}\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}$ is conserved Huo ; HTC . Some methods, such as Green’s function method HTC ; Gera and complete basis expansion method Huo ; Wierzchucka , can be applied to obtain the its eigenstates and eigenenergies. In this subsection, we adopt the Green’s function method to analytically get the eigenstates of the TC model, and its dynamics are subsequently studied, where the presence of disorder gave rise to a variety of complex behaviors. For the HTC model, we utilize the numerically accurate Davydov-Ansatz method, which is confirmed by comparing with the standard basis-set method HTC , to investigate the evolution of photon dynamics.
Below we consider the special case of $N_{\rm ex}=1$ because the manifold of single-photon excitation is the most relevant for applications. We note that our methodologies presented in this Perspective can be readily extended to the manifolds of multi-photon excitations. Thus, the state vectors span the closed ($N$+1)-dimensional Hilbert space with the basis set $\{|1_{c},{\bf 0}_{\rm qu}\rangle,|0_{c},1^{1}_{\rm qu}\rangle,\cdots,|0_{c},1^{n}_{\rm qu}\rangle,\cdots,|0_{c},1^{N}_{\rm qu}\rangle\}$. Here we assume that the cavity mode is initially excited to the state $|1_{c},{\bf 0}_{\rm qu}\rangle$, while the qubits are in their ground states. The Green’s function $\mathcal{\hat{G}}(z)$ of the TC model in the matrix form can be written as
$\mathcal{\hat{G}}(z)=\begin{bmatrix}{\rm\hat{G}^{-1}_{ph}}(z)&-\bf{g}\\ -{\bf g}^{\rm T}&{\bf\hat{G}}_{{\rm qu}}^{-1}(z)\end{bmatrix}^{-1},$ (4)
where ${\rm\hat{G}_{ph}}(z)=(z-\omega_{c})^{-1}$ and $[{\bf\hat{G}}_{{\rm qu}}(z)]_{m,n}=\delta_{m,n}(z-\omega_{n})^{-1}$ are, respectively, free photonic and qubit Green’s functions. The vector ${\bf g}=(g_{1},g_{2},\cdots,g_{N})$ denotes the coupling between the photon and the qubits. The photonic Green’s function is explicitly expressed as
$\mathcal{\hat{G}_{\rm ph}}(z)=\mathcal{\hat{G}}_{1_{c},1_{c}}(z)=\frac{1}{{\rm\hat{G}^{-1}_{ph}}(z)-{\bf g}{\bf\hat{G}}_{{\rm qu}}(z){\bf g}^{\rm T}},$ (5)
with the self-energy
${\bf g}{\bf\hat{G}}_{{\rm qu}}(z){\bf g}^{\rm T}=\sum_{n=1}^{N}\frac{g_{n}^{2}}{(z-\omega_{n})}.$ (6)
The photonic Green’s function $\mathcal{\hat{G}}_{1_{c},1_{c}}(z)$ is directly related to a number of physical quantities, such as the transmission coefficient of the waveguide, population evolution of the photonic mode, and the emission or absorption spectrum. The off-diagonal matrix elements of $\mathcal{\hat{G}}(z)$ are given by the expression
$\mathcal{\hat{G}}_{{\bf 1_{\rm qu}},1_{c}}^{\rm T}(z)=\mathcal{\hat{G}}_{1_{c},{\bf 1}_{\rm qu}}(z)=\frac{{\bf g}{\bf\hat{G}}_{{\rm qu}}(z)}{{\rm\hat{G}^{-1}_{ph}}(z)-{\bf g}{\bf\hat{G}}_{{\rm qu}}(z){\bf g}^{\rm T}},$ (7)
which can be used for the description of correlations between the photon and the qubits. The reader is referred to Appendix A for additional details on the analytic solution of the TC model based on the Green’s function method.
Figure 2: (a) The photon-mode population dynamics evaluated for four diagonal disorder strengths indicated in the legend (the TC model with $N=20$ qubits). (b) Absorption spectra of the TC model for three diagonal disorder strengths for a system $N=100$ qubits. The other parameters are chosen as $\omega_{c}=1\leavevmode\nobreak\ \rm eV$, $\omega_{\rm R}=0.1\leavevmode\nobreak\ \rm eV$, $\omega_{0}=1\leavevmode\nobreak\ \rm eV$. 100 realizations of disorder are taken for obtaining converged results. (c) The photon-mode population dynamics of the HTC model with $\lambda=0.5$ for three diagonal disorder strengths indicated in the legend. The central energy of the phonon band, $\omega_{k0}=0.124\leavevmode\nobreak\ \rm eV$, is adopted. The remaining parameters are the same as in panel (a). (d) Absorption spectra calculated of the HTC model with $\lambda=0.5$ for three values of energy disorder indicated in the legend. Adapted from HTC . Copyright AIP Publishing.
Using Eq. (28), we study a disordered TC model with $N=20$ qubits. The photon-mode population dynamics for increasing disorder strengths ($W=0,0.1,0.2,0.3\leavevmode\nobreak\ \rm eV$) are plotted in Fig. 2(a) without finite lifetimes. The populations exhibit oscillatory behavior, and their median values decrease with static disorder. This can be understood as follows. Static disorder generates higher spread of the systems eigenenergies and richer energy spectrum. Hence more excitation remains localized in the qubits, because the back energy transfer to the photonic mode becomes hindered. Diagonal disorder thus promotes the energy transfer to the qubits. Furthermore, the periods of oscillations of the photonic-mode populations decrease with static disorder. The origin of this phenomenon is clarified by Eq. (22) for the energy-gap frequency. Strictly speaking, Eq. (22) is correct without disorder only. However, it remains qualitatively valid for non-too-high disorder if we average it over the relevant static-disorder distribution. Since $\omega_{0}$ enters the energy gap $\epsilon$ through $(\omega_{0}-\omega_{c})^{2}$, this averaging can only increase the energy gap ($\epsilon_{\rm av}>\epsilon$) and therefore decrease the period of the energy-gap oscillations. Note also that the short-time population dynamics is only slightly affected by diagonal disorder, because the difference between $\epsilon_{\rm av}$ and $\epsilon$ is relatively minor, and several periods are required to produce a substantial phase difference generated by the two frequencies.
[…]
Proper inclusion of vibrational modes is essential for a comprehensive and realistic depiction of organic molecular cavity QED systems. By incorporating phonon modes and qubit-phonon couplings into Hamiltonian (1), we obtain Hamiltonian (II) of the HTC model. Since the HTC Hamiltonian commutes again with the excitation number operator $\hat{N}_{\rm ex}=\hat{a}^{\dagger}\hat{a}+\sum_{n=1}^{N}\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}$, we can work in the lowest-energy excitation manifold and set $N_{\rm ex}=1$. To handle the newly introduced vibrational DOFs and to accurately solve the multidimensional time-dependent Schrödinger equation governed by the HTC Hamiltonian, we adopt the time-dependent variational method with a trial state from the hierarchy of Davydov Ansätze  25 ; 26 ; 27 ; 28 ; 29 ; yaoyao ; 30 ; skw20a . The multiple Davydov $\mathrm{D}_{2}$ Ansatz (also known as multi-$\rm D_{2}$ Ansatz) with multiplicity $M$ can be written as
$|{\rm D}_{2}^{M}(t)\rangle=\sum_{n}|n\rangle\sum_{m=1}^{M}A_{mn}(t)e^{\left(\sum_{k}f_{mk}(t)\hat{b}_{k}^{\dagger}-\mathrm{H.c.}\right)}|0\rangle_{\mathrm{ph}},$ (8)
where $|n\rangle$ stands for the ($N$+1)-dimensional basis state, $|0\rangle_{\mathrm{ph}}$ is the phonon vacuum state, $\rm H.c.$ represents the Hermitian conjugate, $A_{mn}(t)$ are the amplitudes of the excited states, and $f_{mk}(t)$ are the harmonic mode displacements for the $m$th coherent state and the $k$th mode. The multi-D2222 Ansatz includes a superposition of multiple coherent states for each boson mode, which helps deliver accurate wave packet motion on the potential energy surfaces, thus leading to the convergent numerically accurate results for a sufficiently large multiplicity $M$. The approach employing the multi-D2 Ansatz belongs to a large family of methods which use time-dependent Gaussian basis functions to accurately solve multidimensional time-dependent Schrödinger equations. The reader is referred to Appendix B for derivation details on the time evolution of the multi-D2 Ansatz for the HTC model as well as calculations of related physical observables.
The time-dependent photon-mode populations of the HTC model are plotted in Fig. 2(c) for different diagonal disorder strengths indicated in the legend with the qubit-phonon coupling coefficient $\lambda=0.5$. A linear phonon dispersion is assumed to be of the form
$\omega_{k}=\omega_{k0}[1+\Omega(\frac{2|k|}{\pi}-1)],$ (9)
where $\omega_{k0}$ denotes the central energy of the phonon band, $\Omega\in[0,1]$ is the band width, and the momentum is set to be $k=2\pi{l}/N$ with $(l=-\frac{N}{2}+1,\cdots,\frac{N}{2})$. Here the band width of the phonon spectrum is set to $\Omega=0.5$. The populations do not show any signatures of coherent vibrational effects and exhibit oscillations caused by the coupling of the photonic and qubit modes. The qubit-phonon coupling significantly reduces amplitudes of these oscillations. Furthermore, the photon-mode populations gradually approach steady-state values that decrease with $W$, indicating that stronger qubit-phonon coupling facilitates photon-to-phonon mode population transfer. It is apparent that the qubit-phonon coupling plays a role similar to that of the diagonal disorder, in agreement with Ref. GB21 which shows that static disorder can be handled on equal footing with the dynamic disorder caused by vibrational modes. From figs. 2(c), it reveals that strong diagonal disorder lead to the localization of qubit excitations.
The influence of the energy disorder on the absorption spectra for the HTC model with $\lambda=0.5$ is elucidated by Fig. 2(d). Qualitatively, disorder makes dark vibronic polariton states partially bright, owing to incomplete destructive interference. This produces new spectral features in the absorption spectra which are shown in Fig. 2(d) for various diagonal disorder strengths. The spectral plateau sandwiched between the two peaks gradually rises, and the polariton splitting widens with $W$. This is attributed to the disorder-induced dipole moment of the dark vibrational polariton. For $W=0.1\leavevmode\nobreak\ \rm eV$, the mirror symmetry with respect to $\omega=1$ eV is broken, i.e., so-called the polaron-decoupling effect. The electronic and vibrational degrees of freedom contributing to the LP state become separable, which explains relative insensitivity of this peak to the qubit-phonon coupling. On the contrary, the UP peak is significantly broadened by increasing $\lambda$ and exhibits new spectral features. The LP peak largely retains its shape, while the amplitude, position, and width of the UP peak change markedly. Takahashi et al. utilized the HTC model for the interpretation of their steady-state and time-resolved spectroscopic measurements. They demonstrated that when the cavity photon energy is altered, the rate of SF resonantly excited into the LP state is indeed modulated. This finding implies that polaron decoupling is crucial in nonadiabatic dynamics Takahashi . In addition, the extent of polaron decoupling is highly sensitive to the ratio of Rabi splitting to the reorganization energy. Once this ratio surpasses a critical value, complete decoupling occurs, which in turn may modify the chemical reaction rates Zeb18 .
II.2 Dynamics of the non-Hermitian HTC model
Hamiltonian (II) is Hermitian, but in a realistic cavity system photon loss is inevitable due to the following factors: geometric deflection, diffraction, incomplete reflection of cavity mirrors, and non-activated absorption and scattering within materials, which greatly affects the behaviors of cavity assisted QED. To take into account various dissipative effects that are not included in the reservoir Hamiltonian $\hat{H}_{\rm R}$, such as cavity loss, a non-Hermitian term zhang24 ,
$\hat{H}_{\rm nH}=-i\kappa|1_{c},{\bf 0}_{\rm qu}\rangle\langle 1_{c},{\bf 0}_{\rm qu}|,$ (10)
can be added to Eq. (II), where $\kappa$ is the loss rate. The effective Hamiltonian with the non-Hermitian term neglects, in comparison with the Lindblad master equation, the fluctuation term $2\kappa\hat{a}\rho\hat{a}^{\dagger}$. A more sophisticated scheme includes the coupling between the cavity mode and the environmental modes (namely, the Gardiner-Collett interaction Hamiltonian), characterized by a continuum spectral density Huo . This Hamiltonian can be rigorously derived from the QED first-principles. It has been used to investigate polariton quantum dynamics in a dissipative cavity. By constructing an effective non-Hermitian Hamiltonian or employing spectral discretization techniques, the time-dependent variation method with the mD2 Ansatz can be applied to account for dissipation scenarios.
For a nonzero cavity loss rate, both the upper and lower polariton states experience a depletion of their population, transferring it to the ground state. This loss of the excited state population typically diminishes the capacity of the system to undergo reactions on excited surfaces. However, cavity loss may also enhance the rate of photochemical reactions Torres . Thus, cavity loss can be viewed as an additional adjustable element for controlling photochemical polaritonic reactivity.
Figure 3: The photon-mode population dynamics with the finite lifetimes $\kappa=0\rm eV,0.002\rm eV,0.004\rm eV,$ and 0.006 eV for $\lambda=0.1$ (a) and $\lambda=0.4$ (b). The other parameters are set by $\omega_{R}=0.1\rm eV$, $N=10$, $\omega_{c}=\omega_{n}=1\rm eV$, and $\Omega=0.5$. Adapted from zhang24 . Copyright AIP Publishing.
The photon population evolutions $\langle D_{2}^{M}(t)|\hat{a}^{\dagger}\hat{a}|D_{2}^{M}(t)\rangle$ for different cavity loss parameters $\kappa$ and qubit-phonon couplings $\lambda$ are plotted in Fig. 3. As the cavity loss increases, the photon population gradually decreases. Due to the finite lifetimes ($\kappa\neq 0$), the population decays to zero in the long-time limit. The bright-state energy gap for the TC model is $\epsilon=\sqrt{(\omega_{c}-\omega_{0})^{2}+4\omega_{R}^{2}}=2\omega_{R}$, which determines the oscillation period $2\pi/2\omega_{R}\sim 20.7\leavevmode\nobreak\ \rm fs$ of the photon population. As a comparison, for the case of weaker qubit-phonon coupling in the HTC model as shown in Fig. 3(a), we can still observe the regular 20.7 fs periodic oscillations. For $\lambda=0.4$, the phonon-qubit coupling significantly reduces the amplitudes of these oscillations, as shown in Fig. 3(b). Stronger qubit-phonon coupling quenches Rabi oscillations, due to the increasing number of quantum states participating in the HTC dynamics. It is worth mentioning that a partial recurrence of photon population occurs after 150 fs. We attribute it to the limited number of discrete phonon modes in the model. In addition, the period of Rabi oscillations has a small increment for a stronger phonon-qubit coupling strength in Fig. 3(b). Hence, the renormalized Rabi frequency is slightly reduced by the qubit-phonon coupling.
II.3 Finite-temperature dynamics of the HTC model
Figure 4: Evolution of photon population for different qubit-phonon diagonal coupling strengths and temperatures. $\omega_{k0}=0.0124\mathrm{eV}$, $\omega_{\mathrm{c}}=1\mathrm{eV}$, $\omega_{n}=1\mathrm{eV}$ (a)$\lambda=1$, $T=100$K, $T=200$K, $T=300$K;(b)$\lambda=1.5$, $T=100$K, $T=200$K,$T=300$K; (c)$\lambda=2$, $T=100$K, $T=200$K, $T=300$K;(d)$T=300K$, $\lambda=1$, $\lambda=1.5$, $\lambda=2$. Absorption spectra at different temperatures and qubit-phonon coupling strengths $\omega_{k0}=0.0124\mathrm{eV}$. (e)$\lambda=2$, $T=100\leavevmode\nobreak\ \rm K$, $T=200\leavevmode\nobreak\ \rm K$, $T=300\leavevmode\nobreak\ \rm K$. (f)$T=300\leavevmode\nobreak\ \rm K$, $\lambda=1$, $\lambda=1.5$, $\lambda=2$. Adapted from HouE24 . Copyright AIP Publishing.
[…]
In this subsection, we discuss the finite-temperature dynamics of the HTC model, Hamiltonian (3), in which only the diagonal qubit-phonon coupling is considered. Time evolutions of the photon populations for different coupling strengths and temperatures is show in Fig. 4. The temperature effect on the photonic dynamics is relatively weaker for smaller $\lambda$, as shown in Figs. 4(a)-(c). Coherent oscillation of the photon populations is gradually weakened with the increase of the temperature. Here the temperature effect is related to the qubit-phonon coupling. It is much clear if we roughly express the dressed qubit-phonon coupling strength as $\sum_{k}\lambda\cosh\theta_{k}+\sum_{k}\lambda\sinh\theta_{k}$, which increases with temperature, from 1 at $\rm T=0$ to infinity as $\rm T\to\infty$. Obviously, the qubit-phonon coupling $\lambda$ plays the role of an amplification parameter. Besides, the expression of dressed qubit-phonon coupling also tells us that the qubit-phonon coupling acts similarly to temperature, which is further verified by Figs. 4(c) and (d). Hence, significance of temperature effects in the HTC model is determined by two conditions: (1) $k_{B}T$ is comparable with the phonon frequencies $\omega_{k}$. (2) $\lambda$ is not very weak HouE24 .
The HTC absorption spectra without detuning are plotted in Fig. 4. There are two peaks, i.e., the LP and UP peaks, which are approximately located at $0.9\rm eV$ and $1.1\rm eV$, respectively. They are mainly attributed to the bright states with energies $E_{\pm}=1/2(\omega_{c}+\omega_{0})\pm 1/2\sqrt{(\omega_{c}-\omega_{0})^{2}+4\omega_{R}^{2})}$. The separation of the spectral peaks reveals the energy gap $\epsilon\sim 2\omega_{R}=0.2\leavevmode\nobreak\ \rm eV$. As the temperature increases, the peak intensities are weaken, and the separation of the spectral peaks slightly increase due to enhanced dynamics disorder, as shown in Fig. 4(e). It is worth mentioning that the middle portion of the absorption spectrum between the peaks gradually increases in height with the increasing temperature due to the peak broadening. As expected, similar spectral behaviors are exhibited in Fig. 4(f) for different qubit-phonon coupling strengths. Since smaller Huang-Rhys factors are taken, symmetric two peaks are found even though the polaron-decoupling threshold $12N\omega_{R}/\lambda^{2}\omega_{k}=$ 1 of the LP state is satisfied.
Our results are consistent with the observation by Laitz et al. that exciton-phonon scattering decreases as the temperature drops. This decrease leads to an alteration of the bottleneck effect in cavity-controlled perovskite films. Simultaneously, by adjusting the detuning of the microcavity, the optical properties of the perovskite films, such as photoluminescence quantum yield, are greatly changed. Obviously, the temperature control and the cavity detuning can be regarded as the new “knobs” to regulate the photochemical process Laitz2023 . However, the open question is how temperature-dependent polariton relaxation mechanisms affect the rate and product selectivity of polariton-mediated chemical reactions.
III CI-mediated SF in optical cavities
Our aim in this section is to highlight the status quo of the studies into the CI-driven cavity-mediated SF from the point of view of theoreticians. To enhance the readability of this work, a bottom-up approach is adopted. Specifically, we focus on a single representative SF species, rubrene dimers (RDs), and conduct an in-depth analysis of a well-characterized QED system (two RDs in a single-mode cavity). Furthermore, we explicitly monitor the first (ultrafast) leg of the SF process, in which the singlet exciton produces a correlated triplet exciton pair. The process that follows transforming the correlated triplet pair into free triplet states occurs on a much longer timescale and is not considered. We start by reviewing the fundamental processes that shape SF mechanisms and highlighting their spectroscopic signatures. Subsequently, we place these findings within the broader context of theoretical investigations into PPP molecular systems and materials, outlining open questions and computational challenges on the horizon. All simulations here have been performed using time-dependent variation with the multi-D2 Ansatz (cf. Appendix C), a well-established many-body wave-function-based variational integrator of quantum dynamics employing Gaussian coherent states zhao2 ; zhao1 . Provided that the multiplicity of the multi-D2 Ansatz is sufficiently high (this is so for all numerical illustrations of the present work), the trial state can be considered numerically accurate, effectively capturing and faithfully reproducing the dynamics of many-body quantum systems with multiple DOFs.
A molecular (hence the subscript M) system consisting of two rubrene dimers coupled to a cavity photonic mode described by the Hamiltonian
$H=\sum_{j=1,2}H_{{\rm M}}^{(j)}+H_{{\rm C}}+H_{{\rm CM}}^{(j)}.$ (11)
Here $H_{{\rm M}}^{(j)}$ refers to the $j$th pristine (without cavity) dimer, $H_{{\rm C}}$ specifies the cavity field, and $H_{{\rm CM}}^{(j)}$ is responsible for the dimer-cavity photonic coupling. In the diabatic representation, each dimer is modeled as a five-electronic-state two-vibrational-mode system sun2 :
$H_{{\rm M}}^{(j)}$ $=$ $\sum_{k={\rm g}^{(j)},{\rm S}_{1}^{(j)},{\rm TT}^{(j)},{\rm S}_{n}^{(j)},{\rm TT}_{n}^{(j)}}|k^{(j)}\rangle(\varepsilon_{k}+h_{k}^{(j)})\langle k^{(j)}|$ (12)
$+$ $(|{\rm S}_{1}^{(j)}\rangle\langle{\rm TT}^{(j)}|+|{\rm TT}^{(j)}\rangle\langle{\rm S}_{1}^{(j)}|)\lambda Q_{cu}^{(j)}.$
Here $|{\rm g}^{(j)}\rangle$ denotes the electronic ground state, $|{\rm S}_{1}^{(j)}\rangle$ is the singlet electronic state coupled to the triplet-pair state $|{\rm TT}^{(j)}\rangle$, while $|{\rm S}_{n}^{(j)}\rangle$ and $|{\rm TT}_{n}^{(j)}\rangle$ are higher-lying singlet and triplet states. $\lambda$ is the coupling constant specifying the $|{\rm S}_{1}\rangle$-$|{\rm TT}\rangle$ CI responsible for the SF, $\varepsilon_{k}$ are the electronic vertical excitation energies, and $h_{k}^{(j)}$ are the vibrational Hamiltonians involving a single interstate coupling mode (subscript $cu$) and a single primary tuning mode (subscript $tu$):
$h_{k}^{(j)}=h_{g}^{(j)}+\kappa_{k}Q_{tu}^{(j)},$
$h_{g}=\frac{1}{2}\sum_{\alpha=cu,tu}\omega_{\alpha}([P_{\alpha}^{(j)}]^{2}+[Q_{\alpha}^{(j)}]^{2})$
($\hbar=1$). Here $Q_{\alpha}^{(j)}$ and $P_{\alpha}^{(j)}$ are dimensionless coordinates and momenta of the interstate coupling ($\alpha=cu$) and tuning ($\alpha=tu$) modes with frequencies $\omega_{\alpha}$, and $\kappa_{k}$ are the linear intrastate electron-vibrational coupling constants. The effective model of the CI-mediated SF process has been developed in Ref. Miyata and further extended in Refs. Sun6 ; skw . In this model, the tuning mode $Q_{tu}$ is the dimensionless reaction coordinate of the SF process, while $Q_{cu}$ is an antisymmetric interstate coupling mode which couples the states $|{\rm S}_{1}^{(j)}\rangle$ and $|{\rm TT}^{(j)}\rangle$. This model is augmented with the higher-lying excited electronic states $|{\rm TT}_{n}^{(j)}\rangle$ and $|{\rm S}_{n}^{(j)}\rangle$.
The cavity-mode Hamiltonian reads
$H_{{\rm C}}=\omega_{{\rm C}}C^{\dagger}C$
where $C^{\dagger}$ ($C$) is creation (annihilation) operator of the cavity photon with frequency $\omega_{{\rm C}}$. Within the RWA and the dipole approximation, interaction of the dimer with the electric field of the cavity mode is specified by the Hamiltonian
$H_{{\rm CM}}^{(j)}$ $=$ $\frac{\Omega}{2}(CX^{\dagger}+C^{\dagger}X),$ (13)
$X^{\dagger}$ $=$ $|{\rm S}_{1}^{(j)}\rangle\langle{\rm g}^{(j)}|+\eta_{S}|{\rm S}_{n}^{(j)}\rangle\langle{\rm S}_{1}^{(j)}|+\eta_{T}|{\rm TT}_{n}^{(j)}\rangle\langle{\rm TT}^{(j)}|$
where $\Omega$ is the vacuum Rabi frequency for a single molecular emitter. For $N$ identical molecules in the cavity, the Rabi splitting in the Franck-Condon region can be approximated by the formula
$\Delta_{R}=\sqrt{(\omega_{C}-\varepsilon_{S_{1}})^{2}+N\Omega^{2}}$ (14)
derived in the TC model (see, e.g., Ref. HTC and discussion therein). The state ${\rm S}_{1}^{(j)}$ is optically bright from the ground state while the state ${\rm TT}^{(j)}$ is optically dark, so only the ${\rm S}_{1}^{(j)}$ state is coupled to the cavity in Eq. (13). Cavity-induced transitions between the lower-lying and upper-lying states ${\rm S}_{1}^{(j)}$-${\rm S}_{n}^{(j)}$ and ${\rm TT}^{(j)}$-${\rm TT}_{n}^{(j)}$ are governed by the parameters $\eta_{S}$ and $\eta_{T}$. The total Hamiltonian $H$ commutes with the number operator $N_{{\rm ex}}=C^{\dagger}C+\sum_{j,k}|k^{(j)}\rangle\langle k^{(j)}|$ and the total number of excitations is therefore conserved. The present SF model is constructed to investigate the nonlinear spectroscopic signatures in Section III.3, in which the ground-state bleach (GSB), the stimulated emission (SE), and the excited-state absorption (ESA) are included. It is worth mentioning that Hamiltonian (11) is equivalent to the HTC model in the special case of $\lambda=0$.
To explore the optimal parameters of a highly efficient cavity-controlled SF process, the higher-lying states excitation in Hamiltonian (12) can be neglected in the simulations for brevity. Thus, every rubrene molecule mediated by a CI could be modeled as a three-electronic-state (a ground state $\rm g$, an excited singlet $\rm S_{1}$ and a correlated triplet pair state ($\rm{}^{1}(TT)$sun1 . Furthermore, in order to achieve a larger tunable model parameter space, for example, the higher average photon number state, the non-RWA terms should be taken into account in the cavity-molecule coupling Hamiltonian, which reads
$H_{\rm CM}^{(j)}$ $=$ $\frac{\Omega^{(j)}}{2}(|{\rm g}^{(j)}\rangle\langle{\rm S_{1}}^{(j)}|+{\rm H.c.})(C^{\dagger}+C).$ (15)
The non-RWA terms are particularly significant in physical experiments involving strong and ultrastrong coupling Ann21 , short-time quantum dynamics Zheng08 , high-precision spectral analysis Li13 , as well as interactions between ultrashort laser pulses and matter Lu16 . The non-RWA Hamiltonian (15) is adopted to study the SF dynamics with different average photon numbers in Section III.2.
Table 1: Summary of bare SF material parameters and their optimal gain rates of TT yield $\tilde{\epsilon}_{\rm TT}/\epsilon_{\rm TT}$ in the cavity. where $\Delta\rm G$ is given by $\Delta\rm G=\omega_{\rm TT}-\omega_{\rm S_{1}}$, and $\epsilon_{\rm TT}$ ($\tilde{\epsilon}_{\rm TT}$) represents bared (polariton-assisted) TT yield, respectively.
$\rm Molecule$ $\Delta\rm G(meV)$ $\epsilon_{\rm TT}(\%)$ $\tilde{\epsilon}_{\rm TT}/\epsilon_{\rm TT}$ Ref.
$\rm Tetracene$ $150$ 0.01 $\sim 16$ Mart
$\rm Pentacene$ $-110$ 100 $\sim 1.4$ Mart
$\rm Hexacene$ $-630$ 48 $\sim 3.3$ Mart
$\rm Rubrene$ $50$ 17,40,>99.917,40,>99.917 , 40 , > 99.9 $\sim 1.3$ Takahashi ; Maslennikov ; Malin13
The parameters characterizing several popular bare and cavity-manipulated SF materials are listed in Table 1. As follows from this Table, incorporation of these materials in microcavities increases SF yield. Hence integration of QED-cavities into the design of photovoltaic devices offers a notable potential to enhance singlet-to-triplet conversion efficiency and circumvent the Shockley-Queisser limit SQ ; Markvart22 . However, the fulfilment of this program requires understanding of the microscopic mechanisms of the cavity-enhanced SF. In this context, it is essential that the majority of the SF materials (for example, pentacene Musser ; HGD20 , tetracene Herbert17 ; Xiao20 , and rubrene Miyata ) exhibit conical-intersection (CI) driven SF.
CIs are points of degeneracy in the potential energy surfaces (PESs) of isolated polyatomic molecules. Manifolds of these points of degeneracy, known as CI seams, undermine the Born-Oppenheimer approximation and open up effective and fast channels for the population and charge transfer ConicalIntersections . In PPP systems, the situation is complicated by the fact that CIs may be either intrinsic or cavity-induced Mukamel16 ; Kowalewski17 . In the strong coupling regime, an intrinsic molecular CI splits into a pair of polaritonic CIs Gu2 ; sun1 . Furthermore, strong coupling to the cavity mode may produce so-called light-induced CIs Cederbaum12 ; Cederbaum13 , which occur even in diatomic molecules possessing no intrinsic CIs Cederbaum08 ; Csehi . These light-induced CIs may be caused by the presence of additional (for example, rotational/orientational) DOFs. Hence the coupling to the cavity creates multidimensional PES landscapes comprising cascades of intrinsic and cavity-induced/modified CIs Sun6 ; Gu3 ; Gu2 ; sun1 ; Cho ; Gu ; Gu1 ; sun2 . On the one hand, this enormously complicates the situation and imposes high requirements on the numerical efficiency and accuracy of the theoretical methods used for the simulation of dynamic and spectroscopic responses of these systems. On the other hand, the cavity-induced/modified CIs create additional pathways and shortcuts for the population transfer, which can be used for engineering/optimizing the SF process Gu3 ; sun1 . Yet, the polaritonic CIs in the strong-coupling regime may act as double-edged sword. For example, Gu et al. uncovered that SF can be significantly suppressed, since polaritonic CIs are shifted away from the Franck-Condon region in pentacene dimers Gu2 . Hence enhancing the SF process in cavities remains a topic of contention.
III.1 Polaritonic CI landscapes
In rubrene, the energy of the TT state, $\rm E_{TT}$, is slightly higher than the energy of the singlet state, $\rm E_{S_{1}}$. Hence SF – without thermal activation Miyata or excitation to higher-lying electronic states $\rm E_{S_{n}}$ Malin12 – is rather inefficient. Furthermore, the usual mechanism of polaritonic enhancement of SF through the cavity-finetuned resonance of the LP and TT states, is ineffective in rubrene, too Ribeiro2018 ; Takahashi ; Mart . Hence rubrene is a suitable system for exploring novel cavity-induced mechanisms of the SF enhancement.
Figure 5: Cuts of the adiabatic PESs along the tuning mode ($Q^{(j)}_{c}=0$) for one (left column; panels a, b) and two (right column; panels c, d) RDs in the cavity for the first few polaritonic manifolds $N$. (a) No coupling to the cavity mode ($\Omega=0$), the bare CI is located in the Franck-Condon region at $Q^{(1)}_{t}=0.07,E=2.256\leavevmode\nobreak\ \rm eV$. (b, c, d) Strong coupling to the cavity mode ($\Omega=0.2\leavevmode\nobreak\ \rm eV$). The fraction of the TT states ($W_{\rm TT}$ of Eq. (18)) is indicated through the line color coding of PESs (a, c), while fraction of the photonic states ($W_{\rm ph}$ of Eq. (17)), is indicated through the line color coding of the PESs (b, d). The dash lines in panel (b) show positions of the polaritonic CIs. Adapted from sun1 . Copyright American Chemical Society.
Following Ref. sun1 , we consider a microcavity with one or two RDs placed into it. To focus on the net effect of the cavity-induced control mechanisms, the loss of photons in the cavity is neglected. This approximation is reasonable. For example, ultrahigh-Q PhCnB cavities with a quality factor of approximately $\rm Q\sim 750,000$ have been fabricated through the utilization of a five-hole taper design Deotare . Moreover, the nanobeam cavity has been optimized to ensure a quality factor of 1700 and a transmission efficiency of $90\%$. When integrated with the bowtie, the hybrid photonic-plasmonic cavity attains a quality factor of 800 and a transmission rate of $20\%$ Conteduca . Thus, the timescale of the studied SF processes (hundreds of femtoseconds) is short in comparison with the lifetimes of such cavities. In this setup, the $j$th RD participating in the CI-mediated SF is modeled as a system with three electronic states (the ground state ${\rm S}_{0}^{(j)}$, the singlet excited state ${\rm S}_{1}^{(j)}$, and the correlated triplet pair state ${\rm TT}^{(j)}$) and two vibrational modes (the tuning mode with a frequency $\Omega_{t}=0.186$ eV and coordinate $Q^{(j)}_{t}$, and the coupling mode with a frequency $\Omega_{c}=0.0154$ eV and coordinate $Q^{(j)}_{c}$iyata ; sun1 . The electronic energies of the singlet and triplet states are fixed at $E_{{\rm S}_{1}}^{(j)}=2.23$ eV and $E_{{\rm TT}_{1}}^{(j)}=2.28$ eV. The cavity is modelled via a single photonic mode of a frequency $\omega_{c}=2.256$ eV, which is nearly in the electronic resonance with the ${\rm S}_{1}^{(j)}$ and ${\rm TT}_{1}^{(j)}$ states. $\Omega$ determines the coupling of the RDs to the cavity mode. If not explicitly stated otherwise, we set $\Omega=0.2$ eV, which corresponds to the strong cavity-RD coupling regime. To increase space of the adjustable parameters, the RWA in the description of the cavity-matter interaction was not applied.
All levels of the combined RD(s) + cavity system can be classified according to the expectation value of the number operator
$N=\langle\Psi_{n}({\bf Q})|\hat{N}_{\rm ex}|\Psi_{n}({\bf Q})\rangle$ (16)
where $|\Psi_{n}({\bf Q})\rangle$ is the adiabatic eigenfunction of the polaritonic Hamiltonian (since the cavity mode is in electronic resonance with the bright RD state, the RWA is almost exact and the number of excitation is nearly conserved). $N$ specifies the excitation manifold of the polaritonic PESs. For example, $N=0$ corresponds to the collective ground state, $N=1$ yields the singly-excited manifold, etc. To characterize and visualize various states, it is elucidating to introduce several quantifiers. For instance,
$W_{\rm ph}=1-\sum_{\boldsymbol{k}}\sum_{n_{c}\neq 0}|\langle\boldsymbol{k},n_{c}|\Psi_{n}({\bf Q})\rangle|^{2}$ (17)
gives a fraction of the pure photonic excitation in a certain manifold $N$. The sum in Eq. (17) runs over the components $k_{j}={\rm S}^{(j)}_{1},{\rm TT}^{(j)}$ of vector $\boldsymbol{k}$. The number of these components matches the number of RDs in the cavity, and $n_{c}$ is the quantum number specifying the photonic-mode excitation. Similarly, the TT component of the polaritonic state can be defined as
$W_{\rm TT}=\sum_{\boldsymbol{k}\in{\rm TT}}\sum_{n_{c}}|\langle\boldsymbol{k},n_{c}|\Psi_{n}({\bf Q})\rangle|^{2},$ (18)
where the notation $\boldsymbol{k}\in{\rm TT}$ means that at least one of the components $k_{j}$ must contain ${\rm TT}^{(j)}$.
The fractions of the TT states ($W_{\rm TT}$ of Eq. (18)) and photonic states ($W_{\rm ph}$ of Eq. (17)) give a convenient visualization of the polaritonic PESs, which is implemented through the line color coding in Figure 5. Panels (a, b) represent a single RD in the cavity, and show the cut of the polaritonic PESs along the symmetric tuning mode $Q_{t}$ at $Q_{c}=0$. Panel (a) corresponds to a cavity-free RD. Hence, ${\rm S}_{1}^{(1)}$ and ${\rm TT}^{(j)}$ are not dressed with the cavity mode, and only the intrinsic RD CI at $Q_{t}=0.07,E=2.256\leavevmode\nobreak\ \rm eV$ exists in this case. It follows that this CI located in the Franck-Condon region provides the only SF pathway. In the presence of cavity, the PESs change substantially, and cascades of polaritonic CIs appear in Figure 5(b). As the coupling to the cavity increases, the two polaritonic CIs move away from the Franck-Condon point, the distance between the pair of polaritonic CIs increases with $N$, and positions of these CIs lie along the V-shaped dash lines in Figure 5(b). A very similar situation takes place in the pentacene dimer, in which the polaritonic CI is pushed away from the Franck-Condon region, suppressing the SF process Gu2 . However, it is essential that the fraction of the hybrid matter-photon PESs increases with $N$ near the Franck-Condon point. This reversely facilitates the cavity-assisted SF.
The cuts of PESs for two RDs strongly coupled to the cavity mode are shown in Figures 5(c) (highlighting the TT character of the PESs) and (d) (highlighting the photonic character of the PESs). First of all the number of the coupled PESs apparently increases with the number of RDs in the cavity. If for a single RD we have 3 PESs for $N\geq 1$, for two RDs there are 5 PESs for $N=1$, and 9 PESs for $N\geq 2$. The PES with the CIs at $Q^{(1)}_{t}=Q^{(2)}_{t}=0.07$ eV, which corresponds to both RDs in the triplet state, exists for all $N$. With the exception of this PES, the cavity-RD coupling repels polaritonic CIs from the central region, but the effect is not so pronounced as in the case of a single RD.
III.2 SF dynamics Mediated by Polaritonic CIs
Let us explore how the polaritonic PESs coupled through cascades of the cavity-induced CIs affect the SF dynamics. Vendrell suggested the number of photons in the cavity, $N$, can be used as a “control knob”, since stimulated emission induced by a cavity resonantly coupled to a specific molecular reaction coordinate strongly depends on $N$ Vendrell1 . To achieve this, it is worthwhile to study the impact of the initial excitation of the photonic mode on the ensuing singlet-to-triplet population transfer. In our simulations, the cavity mode is assumed to be initially, at $t=0$, in the coherent state
$|\mu_{1}(t)\rangle=\exp(-|\mu_{1}|^{2}/2)\sum_{n_{c}=0}^{\infty}\frac{\mu_{1}^{n_{c}}}{\sqrt{n_{c}!}}|n_{c}\rangle$ (19)
where $\mu_{1}$ is the displacement parameter. For a specific $n_{c}$, the main contribution to the coherent state gives the term with $\mu_{1}(0)=\sqrt{n_{c}}$. Hence we will systematically study how the initial preparation of the RD system in the manifold $N$, where the population is initially placed in the singlet state and the cavity mode is in the coherent state with $\mu_{1}(0)=\sqrt{N}$ affects SF in the regime of strong RD-cavity coupling. This is illustrated by Figure 6 for one (a, c) and two (b, d) RDs in the cavity.
Figure 6: Population dynamics of the triplet state (upper panels) and average photon mode populations $P_{\rm cav}(t)$ (lower panels) for several initial photon-mode pumping numbers $|\mu_{1}(0)|^{2}=N$. Left column (panels a, c): one RD. Right column (panels b, d): two RDs. Adapted from sun1 . Copyright American Chemical Society.
Let us begin with the considerations of the upper panels, which show the triplet populations for different $N$. For a single RD, the triplet population – on the average – increases with $N$ up to $N=6$. For example, the population of $\approx 0.28$ is observed at 300 fs for $N=6$, which is approximately twice larger than the population obtained in the same model without a cavity. As long as these photons remain in the cavity and are available for the next “$\rm S_{0}\to S_{1}\to TT\to charge\leavevmode\nobreak\ transfer$” cycle, the SF efficiency will increase. However, the SF efficiency starts to decrease at $N=7$ and is fully suppressed for $N\geq 8$. This effect can rationalized from the different points of view. On the one hand, it can be viewed as excitation localization on the singlet state ${\rm S}_{1}^{(1)}$ and on the photonic mode, which renders the ${\rm TT}^{(1)}$ effectively decoupled and the SF switched off (see Refs. Localization19 ; Localization20 for engineering of long-lived localized states in the cavity-QED systems). A qualitatively similar localization is realized, for example, in the spin-boson model in the regime of the strong coupling of the spin system to the bosonic bath Wu , even at different system preparations LP23 . Alternatively, the effect can be envisaged as an effective annihilation of the nonadiabatic coupling between the singlet and triplet states Gelin09 . It operates through the combined effect of the cavity field (through Rabi cycling) the system itself (through the intrinsic ${\rm S}_{1}^{(1)}-{\rm TT}^{(1)}$ CI), since the photonic mode does not directly couple the singlet and triplet states. Note also that the population evolution of the triplet states starts not from $t=0$, but from a characteristic time $\tau_{N}$ which increases with $N$. This indicates that the wavepacket needs $\approx\tau_{N}$ to reach the polaritonic CIs.
[…]
Figure 7: Energy levels of two rubrene dimers in the cavity. The superscripts $(1)$ and $(2)$ number the dimers.
Let us now inspect evolutions $P_{\rm cav}(t)$ of the cavity-mode populations vs. the initial photon pumping numbers $N$ ($P_{\rm cav}(0)=N$) which deliver a complementary information. $P_{\rm cav}(t)$ for a single RD is depicted in Fig. 6(c). For $N$ from 1 to 6, the populations exhibit similar behavior: During the first 50 fs, they show several oscillations with a rough period of 22 fs which is related to the tuning mode. Then they reach a quasi-steady-state $\langle P_{\rm cav}\rangle$ value at longer times. It is worth mentioning that $\langle P_{\rm cav}\rangle$ are slightly larger than the initial populations $P_{\rm cav}(0)=N$. It reveals the population redistribution due to the coupling of the cavity mode and the ${\rm S}_{1}^{(1)}$ state. $P_{\rm cav}(t)$ for $N=7$ shows high-amplitude oscillations during the first $\sim 40$ fs, which matches the time $\tau_{7}$ during which the corresponding triplet population in Fig. 6(a) is almost zero. For $N=8$, the high-amplitude oscillations persist over the entire evolution of $P_{\rm cav}(t)$, which is a signature of the localization of the population on the photonic mode and the singlet state, which results in the SF stopping. Interestingly, the period of these fast oscillations reveals the effective Rabi frequency $\Omega_{N}\approx\Omega\sqrt{N}$ ($2\pi/\Omega_{8}\approx 7.3$ fs), which is defined under the assumption that the effective coupling to the cavity scales $\sim\sqrt{N}$.
Fig. 6(d) is a counterpart of Fig. 6(c), but for two RDs coupled to the cavity. For $N$ from 1 to 6, $P_{\rm cav}(t)$ reach their first local minima at $t\approx 17$ fs (dashed line in panel d), which indicates that a fraction of the photonic-mode population is transferred to the doubly excited electronic states of the RDs. Then, $P_{\rm cav}(t)$ exhibit oscillations with a period of $\sim 26$ fs, which cannot be directly related to the model parameters. In general, excitation to $1\leq N\leq 6$ reflects population transfer between photonic mode and excited states of the RDs. For $7\leq N\leq 9$, $P_{\rm cav}(t)$ shows fast Rabi oscillations at the initial stage, which correlate with the absence of the triplet excitation at the same time interval (cf. Fig. 6(b)). For $N=10$, fast Rabi oscillations spread over the entire $P_{\rm cav}(t)$ evolution, which is the signature of the SF switching-off in Fig. 6(b). In the long-time limit, the $P_{\rm cav}(t)$ reach quasi-steady states $\langle P_{\rm cav}\rangle$ the values of which are lower than the initial values $P_{\rm cav}(0)$. This indicates a net population transfer from the photonic mode to the doubly-excited electronic states of RDs.
The above analyses demonstrate that the SF efficiency is primarily determined by two factors. The first factor is the position of the polaritonic CIs (see Figure 5). If the CIs are located in the Franck-Condon region, the subsequent population transfer is facilitated and the higher SF efficiency is achieved. The second factor is the initial wavepacket preparation and the state of the cavity which is controlled by the photonic-mode pumping number $|\mu_{1}(0)|^{2}=N$. This later parameter plays the crucial role, increasing the SF efficiency for small $N$ and fully suppressing SF once $N$ reaches a certain critical value. Similar results are obtained in Ref. Vendrell1 . If the number of photons is higher than the number of the available excited states, the effective cavity-matter coupling may increase as well, leading to slower motion of the nuclear wave packet, which gets partially trapped in the region where the molecular electronic energy gap becomes resonant with the cavity.
Furthermore, collective effects, such as polaron decoupling and collective dark states, have significant impacts on the cooperative polariton dynamics Gu2 . SF dynamical pathways can also be modified by strong photon-matter coupling. For instance, the SF channel via the UP state impedes vibronic losses to the dark-states manifold, which evidently enhances the SF efficiency Vendrell3 .
III.3 Spectroscopic signatures of Polaritonic CIs
As aforementioned, strong coupling of RDs to the cavity mode drastically alters the SF efficiency owing to the cohesive action of the cavity-induced and intrinsic CIs. This processes are imprinted into nonlinear femtosecond spectroscopic responses of these systems. To find spectroscopic signatures of the polaritonic SF occurring in multidimensional PES landscapes, the signals must be accurately simulated and adequately interpreted. For this to achieve, it is essential to realistically model RDs. Indeed, all third-order spectroscopic responses consist of three contributions, the GSB, the SE, and the ESA Darius04 ; Leonas . The GSB and SE reflect the photoinduced processes in the manifolds with $N=0$ (the ground state) and $N=1$ (singly-excited states), while ESA involves also the doubly-excited states, $N=2$. Hence modeling individual RDs as systems containing only singly-excited states ${\rm S}_{1}^{(j)}$ and ${\rm TT}^{(j)}$ is insufficient: we have to include higher-excited states ${\rm S}_{n}^{(j)}$ and ${\rm TT}_{n}^{(j)}$ sun2 . Following Miyata ; sun2 , we set $E_{{\rm S}_{n}}^{(j)}=4.33$ eV and $E_{{\rm TT}_{n}}^{(j)}=4.68$ eV in our simulations. For two RDs in the cavity, we thus obtain a single ground state, 5 singly-excited states and 15 doubly-excited states which are specified in Fig. 7. The states, in turn, can be subdivided into the matter states ($n_{c}=0$, left part of Fig. 7) and hybrid states ($n_{c}>0$, right part of Fig. 7).
Figure 8: Real part of the SE (row (a)), GSB (row (b)), ESA (row (c)) contributions and the total 2DES spectrum (row (d)) of the two RD system for several population times $T_{w}$ indicated in the panels. The spectra are computed in the impulsive limit. Adapted from sun2 . Copyright American Chemical Society.
Here we consider electronic two-dimensional (2D) spectra, because electronic 2D spectroscopy (2DES) is one of the most powerful and information-rich third-order spectroscopic techniques (see Scholes22 and references therein). Qualitatively, 2DES spectra $S(\omega_{\tau},T_{w},\omega_{t})$ can be envisaged as a $T_{w}$-dependent 2D $(\omega_{\tau},\omega_{t})$ maps, which reveal correlations between the states of the system excited at a frequency $\omega_{\tau}$ at time zero and probed at a frequency $\omega_{t}$ at time $T_{w}$. 2DES spectra $S(\omega_{\tau},T_{w},\omega_{t})$ simulated for the two RD system strongly coupled to the cavity mode are presented in Fig. 8 for different population times $T_{w}$. From top to bottom, shown are the SE (row (a)), GSB (row (b)), and ESA (row (c)) contributions, as well as the total signal (row (d)). The reader is referred to Appendix B for more derivation details on the various nonlinear response functions within the framework of the multi-D2 Ansatz.
The SE contribution (row (a)) possesses two pronounced diagonal peaks (DPs) and two cross-peaks (CPs) (similar situation takes place in the purely polaritonic dimer system Bondarev13 ). The DP1 ($\omega_{\tau}=2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$) and DP2 ($\omega_{\tau}=2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) reveal the middle polariton (MP) and the UP bands, respectively. The low-intensity LP band cannot be reliably identified in the SE 2DES spectrum, because 2DES peak intensities are proportional to the product of four relevant transition dipole moments, in contrast with absorption spectra, which are proportional to the product of two transition dipole moments (see the discussion in Ref. Jindra15 ). The CP1 ($\omega_{\tau}=2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$) and CP2 ($\omega_{\tau}=2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) quantify the coupling and polaritonic coherence between the states MP and UP. A weak CP ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) at $T_{w}=16$ fs reveals the phonon sideband. The shape and intensity of the DPs and CPs depend significantly on the waiting time $T_{w}$. The DPs at $T_{w}=0$ are predominately elongated perpendicular to the main diagonal, but their elongation turns into parallel at longer population times. A certain similarity of the DPs at $T_{w}=0$ and 32 fs shows a significant contribution of the polaritonic beating with a period of $\sim 32\leavevmode\nobreak\ {\rm fs}$ revealing the energy difference of $0.132\,{\rm eV}$ between the states MP and UP. The two negative CPs are elongated parallel to the main diagonal. Similar to the DPs, the CPs are also significantly modulated by the MP-UP energy-gap beating. The overall decreases of the intensity of the SE contribution with population time mirrors the CI-mediated population transfer from the bright ${\rm S}_{1}^{(j)}$ states to the dark ${\rm TT}^{(j)}_{1}$ states and reflects thereby the SF dynamics.
The GSB contribution is shown in panels (b) of Fig. 8. It reveals the same pairs of the polaritonic DPs and CPs and, additionally, a pair of weaker polariton-phonon CPs, ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) and ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$). In comparison with the SE peaks, their GSB counterparts look rather static. Nevertheless, the GSB contribution reveals the vibrational wave-packet motion in the electronic ground state, which is clearly manifested in the peak evolutions (see Ref. sun2 ).
The ESA contributions are depicted in panels (c) of Fig. 8. The SE and ESA contributions to 2DES spectra reveal photophysical processes in the manifold $N=1$ skw20a . However, the two contributions employ different reporter states for delivering the same information: the SE employs the electronic ground state $N=0$, while the ESA employs the doubly-excited manifold $N=2$. Hence shapes and evolutions of predominantly positive SE contributions and predominantly negative ESA contributions are quite similar. Nevertheless, positions of the DPs and CPs in the SE and ESA spectra are slightly shifted with respect to each other. At $T_{w}=0\leavevmode\nobreak\ {\rm fs}$, for example, the (nearly) diagonal ESA peaks are located at ($\omega_{\tau}\approx 2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.317\leavevmode\nobreak\ {\rm eV}$) and ($\omega_{\tau}\approx 2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.416\,{\rm eV}$). All ESA peaks are elongated parallel to the $\omega_{t}$ axis, due to a high density of states in the manifold $N=2$. The peak ($\omega_{\tau}\approx 2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.317\leavevmode\nobreak\ {\rm eV}$), which revels transitions from the MP state to the states of the doubly-excited manifold, has the highest intensity. Note also that the relative separation between the DPs and CPs along the $\omega_{t}$ axis changes with $T_{w}$: cf. the spectra at 0 fs with those at 16 fs. This modulation is caused by the superposition of polaritonic and purely vibrational tuning-mode oscillations. This reflects a higher complexity of the ESA contribution, which involves polaritonic states with $N=1$ and 2.
Finally, the total 2DES signal $S(\omega_{\tau},T_{w},\omega_{t})$ is shown in Fig. 8(d). Since positions of the DPs and CPs in the SE, GSB and ESA contributions nearly coincide, the total 2DES spectrum exhibits positive, negative and dispersed peak shapes, the amplitudes of which are determined by subtle cancellation effects in the superposition of positive SE and GSB contributions and negative ESA contribution. The overall structure of the total 2DES spectra is dominated by the quartet of polaritonic DPs and CPs. In addition, a number of polariton-vibrational peaks around $\omega_{\tau}\,\mathrm{or}\,\omega_{t}=$ 2.6 eV is seen at different $T_{w}.$ It is remarkable that these hybrid polariton-vibrational peaks are more pronounced in the total signal than in the SE, GSB, and ESA contributions – owing to the aforementioned cancellation effects. The DPs are positive while the CPs exhibit unusual asymmetry: the upper CP is always positive, but the spectral structures around the lower CP can be negative (see the signals for $T_{w}=0$ and 16 fs).
IV Simulation of polaritonic systems: unsolved problems, challenges, and outlooks
IV.1 Model and Hamiltonian
While the essential features of polariton dynamics can be effectively captured using the HTC Hamiltonian, the study of more practical SF-related photophysical processes has multifaceted significance and application value. The direct ${\rm S}^{(1)}\rightarrow{\rm TT}^{(1)}$ CI-driven SF is one of the possible mechanisms. In specific systems, charger transfer, locally-excited, doubly-excited, and multi-excitonic states facilitate the SF process, determine its pathways, and govern its efficiency MT18 ; MT19 . Furthermore, the coupling between these states can be of different nature, namely a dipolar coupling (which corresponds to the avoided crossing between the corresponding electronic/excitonic PESs) or a vibrational-mode mediated coupling (which corresponds to the CI between the corresponding electronic/excitonic PESs). Obviously, embedding a multi-mode, multi-state system into a microcavity produces a cascade of coupled polaritonic states and PESs governing the SF process. If several SF systems are embedded in one cavity, the situation becomes even more complex. It is essential though that all these polaritonic states directly or indirectly contribute to the coupling between the parent state ${\rm S}^{1}$ and the product state ${\rm TT}^{1}$. Hence, their net effect can be treated by introducing an effective ${\rm S}^{(1)}$-${\rm TT}^{(1)}$ coupling strength. We thus anticipate that the dynamical SF effects described here remain qualitatively valid if a large number of relevant electronic states is included.
A proper characterization of cavity-mediated SF requires a careful selection of the relevant excitonic states and mechanisms of their couplings, notably for a large number $N_{M}$ of SF species embedded in the cavity. In this context, it is appropriate to mention a recent discussion on the possibility to reach a strong-coupling regime in the cavities hosting a large number of species. Liu et al. pinpointed significant challenges in modifying photophysical dynamics in an optical microcavity, necessitating a molecular system with long excited-state lifetimes or architectures that enhance exciton-photon coupling strengths. These requirements are stipulated by the existence of the (quasi)degenerate manifold of dark “reservoir” states (e.g., triplets) as well as the slow conversion from reservoir states to cavity polaritons liu1 . In other words, the results of Liu et al. liu1 as well as those of Takahashi and Watanabe Watanabe20 seem to demonstrate that a good portion of the singlet-state population bypasses the correlated triplet state and goes to the lower dark polaritonic states in the strong-coupling regime. On the other hand, a recent report by the Chergui group Chergui20 does not support this scenario. There appears to be a consensus that this challenge can be addressed through a careful and strategic redesign of the entire cavity QED system. Menon24 .
In excitonic systems, doubly-excited states are usually understood as the states in which two chromophores are simultaneously excited. This interpretation, however, is not always sufficient for SF aggregates, because polyatomic molecules comprising these aggregates possess dense manifolds of higher-lying excited electronic states and the minimal system supporting the SF reaction is not a single chromophore, but a dimer. Usually, higher-lying electronic states do not contribute to the SF per se (see the discussion below though). Yet, optical transitions between the lower-lying and higher-lying states affect dramatically ESA contributions to spectroscopic signals. Hence higher-lying states of individual molecules comprising SF dimers, aggregates, and materials have to be taken into account in any realistic simulation of spectroscopic responses.
Taking into account the multimode (essentially, quasicontinuum) spectral densities of the microcavities Demler20 ; Gerrit22a ; Ribeiro22 and finite cavity-mode lifetimes Huo24 may be essential for a proper description of the polaritonic dynamics in SF systems. Thus, a single continuous-mode description of the system-cavity coupling may be insufficient or even inadequate. However, the vertical excitation energy of a single dimer is around 2.2 eV. In this case, an effective single-mode description can be adopted for multi-mode cavities as only those modes in resonance with this transition would substantially contribute to the polaritonic SF. Similar effects have been explicitly demonstrated, e.g., in Ref. GB24 where the fine (400 modes) description of the photonic spectral density was adopted.
It is crucial that all the additional states and specific inter-state couplings mentioned above can be effectively handled using the multi-D2 Ansatz when required. Moreover, the methodology employing multi-D2 Ansatz is efficient for arbitrary time-dependent Hamiltonians, non-Hermitian Hamiltonians, and models of multispecies, multimode bosonic baths zhao2 ; zhao1 . Hence, the multi-D2 Ansatz is a method of choice for handling quasicontinuous spectral densities of the cavity modes, finite lifetimes of the cavity/plasmonic modes, and fine discretization schemes of the cavity modes.
Characterizing all relevant electronic states and parameterizing the SF Hamiltonians using ab initio quantum chemistry methods proves to be highly beneficial. MT18 ; MT19 ; YiZhao21 ; Haibo21 ; Vendrell23 . In addition, ab initio approaches can also be applied to the entire systems of molecules/aggregates including the cavity. For example, quantum mechanics/molecular mechanics (QM/MM) simulations of Refs. Huo ; Gerrit22 ; Gerrit22a treat up to 512 polyatomic chromophores in a cavity. Simulations such as those based on the ab initio QED Rubio23 ; Koch21 , the multiple spawning method Martinez24 , and the cavity Born-Oppenheimer Hartree-Fock Ansatz Kowalewski23 can also provide valuable insights into dynamics of SF reactions in specific systems.
IV.2 Dynamics and spectroscopy
Usually, accurate quantum simulations of up to two CI-carrying molecules/dimers in a cavity are performed Gu1 ; sun1 ; sun2 . Vendrell et al., however, have simulated up to seven pyrazine molecules in a cavity, with each molecule having four modes and three states. They conducted a detailed analysis of the coupling between the cavity and the resonances of $S_{0}/S_{2}$ as well as $S_{1}/S_{2}$ Vendrell . It is of great interest to push the simulations towards higher number of molecules/dimers $N_{M}$ and explore how CIs affect cooperativity, e.g., to check the conventional $\sim\sqrt{N_{M}}$ scaling of the effective system-cavity coupling TAVIS ; Huo22 .
To achieve this aim, a combination of advanced theoretical models and computational techniques appears promising. Few-molecule models with symmetry-optimized Hamiltonians JYZ23a provide a simplified yet effective framework for understanding key interactions. Complementing these models, the multi-D2 Ansatz, as utilized here zhao1 ; zhao2 , offers a powerful variational approach to accurately capture the dynamics of multi-species, multi-modes systems with a large number of DOFs. In particular, the HTC model containing dozens of qubits is readily tackled with the multi-D2 Ansatz technique HTC . Additionally, the multi-configuration time-dependent Hartree (MCTDH) method MT18 ; MT19 ; Vendrell19 ; Vendrell stands out as a robust tool for exploring high-dimensional quantum dynamics, particularly in systems where vibrational and electronic degrees of freedom are strongly coupled. The multi-layer extension of MCTDH has been carefully compared with the multi-$\rm D_{2}$ Ansatz method by Saalfrank and coworkers in a very recent study of vibrational relaxation dynamics at surfaces Fischer . While both rely on the time-dependent variational principle to arrive at an optimal wave function via solving highly nonlinear coupled equations of motion for variational parameters, the mDA method was found to be much more efficient (15,000s on a modern laptop) than the multilayer MCTDH (54,000s on a modern workstation). Furthermore, another potentially useful approach is the so-called time evolving density operator with orthogonal polynomials algorithm (TEDOPA) Prior2010 ; Chin2011 . Through a basis change of the Hamiltonian, the TEDOPA method transforms the environment into a chain of harmonic oscillators with only local nearest-neighbor interactionsGuimares ; Tamascelli . By utilizing Matrix Product States (MPS) for simulating the dynamics of one-dimensional quantum systems, TEDOPA effectively circumvents the otherwise unmanageable scaling associated with system size. Yet, the multi-$\rm D_{2}$ Ansatz method seems to be more flexible; for example, it has been applied very recently to the two-dimensional t-J model shen2024 . Furthermore, the multi-$\rm D_{2}$ Ansatz method can also address systems with long-range or strong interactions, as well as those involving entanglement. These challenges typically require a higher bond dimension in the MPS, leading to a long simulation time with TEDOPA. It is worth mentioning that the tensor-network techniques Vidal18 ; Keeling22 have emerged as an effective approach for handling the computational challenges associated with large-scale quantum systems.
Recently emerging emulations of CIs on quantum simulators CIsimulations23a ; CIsimulations23b ; Schoelkopf23 may also be extended in the future to cavity-CI systems. Unfortunately, the development of QM/MM approaches and numerically accurate fully quantum methods has largely progressed in isolation, with limited integration or cross-validation. This separation creates a significant gap in leveraging the strengths of both methodologies. Fully quantum methods provide high-precision insights into the dynamics of SF systems but are computationally intensive. In contrast, QM/MM methods offer a more computationally feasible framework by combining quantum-level detail with classical mechanics approximations, making them suitable for simulating larger systems or processes. Given these complementary strengths, it is highly desirable to bridge this gap by systematically validating QM/MM methods against fully quantum calculations. Such a validation process would not only enhance the reliability of QM/MM approaches but also enable their broader application in scenarios where fully quantum simulations are computationally prohibitive.
Disorder is of paramount importance to any realistic modeling of CI-driven cavity-mediated SF systems, as defects and disorder in semiconducting molecular materials can give rise to the formation of CIs Chuntonov22 and affect polaritonic transport substantially Schachenmayer22 ; HTC . Temperature adjustment is essential, e.g., for activating CI-mediated SI in rubrene Miyata and for catalysing reactions in polaritonic systems Campos . The multi-D2 Ansatz protocol based on the TFD BG21 was recently applied to explore dynamics of the pristine RD skw ; Sun20 . Subsequently, Hou et al. utilized this approach to obtain the finite-temperature dynamics and the absorption spectra of HTC model HouE24 , as referred to in Sect. II.3. Looking forward, combining the multi-D2 Ansatz with the TFD method, exploration of cavity-controlled SF processes at finite temperatures is anticipated to uncover new mechanisms, including the impact of vibrational polarization on the SF process induced by strong vibrational coupling, the influence of nonlinear coupling terms (such as the Kerr effect) within the cavity, and the role of multi-mode cavities. These factors are expected to provide deeper insights into the interplay between cavity dynamics and SF processes, potentially revealing novel pathways for enhancing efficiency and control. Apart from multi-D2 Ansatz, MCTDH and variants of tensor-network methods are also well-suited for conducting fully quantum simulations of realistic polaritonic CI-driven SF systems, as they enable efficient evaluation of nonlinear spectroscopic signals Santoro23 ; GB_JCTC and – by using the TFD method BG21 and the recipe of Ref. GB21 – effects of finite temperatures and static disorder can also be taken care of in a (relatively) computationally inexpensive manner.
Yet a microscopic and fully quantum description of spectroscopic experiments has to await future developments: The bottleneck is accounting for the large number of system and cavity DOFs and necessity to perform simulations on relatively long (picosecond to microsecond) timescales. An extra challenge is to identify spectroscopic signatures of polaritonic effects in nonlinear spectroscopic signals of polaritonic CI-driven SF systems. Hence, accurate theoretical simulations and interpretations of the existing Takahashi ; Watanabe20 ; Chergui20 and forthcoming experiments are demanding but rewarding.
IV.3 Mechanisms of polaritonic CI-driven SF
There exist several methods that can potentially enhance the SF yield. Conventional practices are based on the adjustment of the cavity mode for achieving resonance and strong coupling of the LP state with the triplet state. Moreover, Vendrell et al. discovered the possibility of using the UP state as an entry door for the cavity-assisted SF pathways in a rigid system, which leads to a much higher yield Vendrell3 .
On the other hand, the geometric-phase effects may also be harnessed for engineering SF in CI-driven polaritonic systems. The controlling mechanisms may be implemented through collective topological polaritonic phases JYZ23 that are akin to those in topological metamaterials Alu23 . It is also worthwhile to explore interfacing of light-induced CIs in cavities Vibok22a ; Vibok22b ; Huo21 with intrinsic CI systems, as their combination may facilitate fine-tuning of the multiple CI landscapes VendrellVibok .
Apart from the aforementioned cavity-related mechanisms, there exist other handles for tweaking SF, which are based on the optimization of the bias voltage Xiong18 , temperature Miyata ; Xiong15 ; Wasielewski21 , carrier frequency BY22 and intensity Asahi16 ; Bardeen18a of the excitation pulse, or even static disorder Chavez ; Disorder24 .
Yet, it is not always clear whether the coupling to the cavity (perhaps in combination with other methods mentioned above) enhances or suppresses the SF yield. It is promising, therefore, to interface the existing protocols for dynamic simulations of polaritonic CI-driven SF systems with quantum control routines Control1 ; Control2 , in order to find a robust pathway to comprehensive optimization of SF in cavities.
V Conclusions
Integrating molecular materials into optical cavities has ushered in a new era of controlling quantum dynamics in complex systems. By coupling molecular excitations with quantized light fields, cavity environments reshape exciton energies, open new reaction pathways, and offer routes to surpass traditional photoconversion limits. This Perspective highlights recent developments in fully quantum simulations to understand CI-driven SF process in an optical cavity. A numerically accurate Davydov trial state, the multi-D2 Ansatz, is employed in a time-dependent variational framework to explore dynamic and spectroscopic responses of the TC model, the HTC model, and the CI-driven SF RD system strongly coupled to a cavity mode. For the disordered TC model, time evolution of the photon population and the optical absorption spectra are also calculated by using the Green’s function method to corroborate numerically accurate results of the multi-D2 Ansatz. The impacts of temperature, static disorder, photon loss, and qubit-phonon coupling on the dynamics and absorption spectra of the HTC model are investigated by combining the multi-D2 Ansatz combined with the TFD method.
Several microscopic SF mechanisms have been uncovered, viz. proximity of the CI-driven polaritonic wave packet to the Franck-Condon region, cavity-induced enchantment/weakening/suppression of SF, polaron/polariton decoupling, localization of the singlet-state population via engineering of the cavity-mode excitation, collective enhancement of SF.
Simulation and inspection of nonlinear femtosecond spectroscopic signals of the polaritonic CI-SF systems provides accurate information on real-time singlet-to-triplet transfer. In particular, 2DES spectra of the RD system are dominated by the quartet of DPs and CPs that reveal the MP and UP bands, while the hybrid polariton-vibrational CPs contain information on the UP sideband. This clear separation of the polaritonic and vibrational peaks is a noticeable feature of the spectra obtained. In addition, simulated 2DES spectra exhibit multiple peak-cancellation effects, and their population-time evolution is accompanied by polariton-vibrational beatings.
The findings emphasize the transformative role of polaritonic states in controlling SF dynamics, highlighting key factors such as cavity mode tuning, Rabi splitting, and electron-phonon interactions. Moreover, the study underscores the importance of integrating advanced simulation techniques, such as the multi-D2 Ansatz, to capture the complex interplay of electronic, vibrational, and photonic states in many-body systems.
Strictly speaking, the level of description of the CI-driven cavity-enhanced SF systems adopted in this perspective and the ensuing effects outlined above pertain to so-called picocavities, in which the strong-coupling regime can be achieved for a few embedded molecules Bhuyan . To validate these predictions, it will be worthwhile to perform ultrafast spectroscopic (perhaps polarization-sensitive) experiments on a series of well characterized SF systems (such as dimers, trimers, and tetramers) in solution and in picocavities. Furthermore, the effects akin to those demonstrated in this Perspective can likely be found in more realistic CI-driven SF systems containing a large number of SF species strongly coupled to multi-mode cavities.
It is concluded that microcavities can open up new avenues for improving efficiency and stability of the CI-mediated SF, and ultrafast nonlinear spectroscopic signals – if interpreted with adequate theoretical support – can provide a real-time guidance on the SF reaction pathways. Yet, polaritonic SF studies have not reached the level of maturity, and many problems await efficient and practical solutions. Looking ahead, the combination of fully quantum simulations, insightful modeling tools, and experimental innovations can guide us toward next-generation SF materials and cavity designs. By addressing these unsolved problems and challenges, researchers have the potential to leverage CIs and cavity architectures for surpassing conventional efficiency limits in solar energy harvesting and paving the way for groundbreaking innovations in next-generation photovoltaics and other optoelectronic applications.
Acknowledgments
The authors gratefully acknowledge the support of the Singapore Ministry of Education Academic Research Fund (Grant No. RG87/20). K. Sun would also like to thank the Natural Science Foundation of Zhejiang Province (Grant No LY18A040005) for partial support. M. F. G. acknowledges support from the National Natural Science Foundation of China (Grant No. 22373028).
Conflict of Interest
The authors of no conflicts to declare.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Green’s function approach to the TC model
Based on Eqs.(4-7), one can investigate the energy spectral structure and the dynamics of TC model. Without static disorder, $\omega_{n}=\omega_{0}$ and $g_{1}=g_{2}=\cdots=g_{N}=\omega_{\rm R}/\sqrt{N}$. Hence the propagators read
$\langle 1_{c},{\bf 0}_{\rm qu}|\hat{U}(t)|1_{c},{\bf 0}_{\rm qu}\rangle$ (20)
[…]
$=$ $\frac{(E_{+}-\omega_{0})e^{-iE_{+}t}-(E_{-}-\omega_{0})e^{-iE_{-}t}}{E_{+}-E_{-}},$
$\langle 0_{c},{1}_{\rm qu}^{n}|\hat{U}(t)|1_{c},{\bf 0}_{\rm qu}\rangle$ (21)
[…]
$=$ $\frac{\omega_{\rm R}/\sqrt{N}e^{-iE_{+}t}-\omega_{\rm R}/\sqrt{N}e^{-iE_{-}t}}{E_{+}-E_{-}}$
(the last lines in Eqs. (20) and (21) are obtained by the residue theory). Here $z=\omega+i0^{+}$ and the eigenenergies of the bright states of the system, $E_{\pm}=1/2(\omega_{c}+\omega_{0})\pm 1/2\sqrt{(\omega_{c}-\omega_{0})^{2}+4\omega_{\rm R}^{2}}$, are obtained from the poles of $\mathcal{\hat{G}}_{1_{c},1_{c}}(\omega+i0^{+})$ corresponding to the eigenmodes strongly coupled to the cavity mode. The bright-state energy gap
$\epsilon=E_{+}-E_{-}=\sqrt{(\omega_{c}-\omega_{0})^{2}+4\omega_{\rm R}^{2}}$ (22)
is therefore formed, and the remaining $N-1$ degenerate dark states, which are decoupled from the cavity mode, are located inside the gap.
Eq. (20) yields the following expression for the population of the photonic mode:
$P_{\rm ph}(t)=|\langle 1_{c},{\bf 0}_{\rm qu}|\hat{U}(t)|1_{c},{\bf 0}_{\rm qu}\rangle|^{2}.$ (23)
Furthermore, the absorption spectrum can be written as
[…]
$=$ $2{\Re}\Big{\{}\frac{E_{+}-\omega_{0}}{E_{+}-E_{-}}\delta(\omega-E_{+})\Big{\}}+$
$2{\Re}\Big{\{}\frac{E_{-}-\omega_{0}}{E_{-}-E_{+}}\delta(\omega-E_{-})\Big{\}}$
(for simplicity, all transition dipole moments are set to 1). To account for dissipation and dephasing caused by the coupling of the system to the environment, we add the non-Hermitian phenomenological lifetime terms $-i\kappa|1_{c},{\bf 0}_{\rm qu}\rangle\langle 1_{c},{\bf 0}_{\rm qu}|-i\Gamma_{n}|0_{c},1^{n}_{\rm qu}\rangle\langle 0_{c},1^{n}_{\rm qu}|$ to the Hamiltonian of Eq. (1). When $\Gamma_{n}=\Gamma$, we can easily extend Eq. (20) by replacing $\omega_{c}$ ($\omega_{0}$) with $\omega_{c}-i\kappa$ ($\omega_{0}-i\Gamma$), respectively. In this case, the absorption spectrum assumes the form
$F(\omega)$ $=$ $\frac{2}{\pi}{\Re}\Big{\{}\frac{E_{+}-\omega_{0}}{E_{+}-E_{-}}\frac{\Im(E_{+})}{[\omega-{\Re}(E_{+})]^{2}+\Im(E_{+})^{2}}\Big{\}}+$
$\frac{2}{\pi}{\Re}\Big{\{}\frac{E_{-}-\omega_{0}}{E_{-}-E_{+}}\frac{\Im(E_{-})}{[\omega-{\Re}(E_{-})]^{2}+\Im(E_{-})^{2}}\Big{\}}.$
Here finite widths of the absorption peaks are determined by the imaginary parts of the eigenenergies. The above results are valid for any number of qubits.
Let us discuss now the influence of static disorder. namely, the usual diagonal disorder in qubit energies $\omega_{n}$. It assume that $M[\omega_{n}^{j}]=\omega_{0}$, where random energies $\omega_{n}^{j}$ are uniformly distributed in $[\omega_{0}-W/2,\omega_{0}+W/2]$ and $W$ are the disorder strengths. Here $M[\cdot]$ denotes the ensemble average over random realizations of $\omega_{n}^{j}$. It is worth mentioning that the analytical results can be obtained if the Gaussian distribution disorder is adopted Gera .
Evidently, disorder removes degeneracy of the eigenenergies. In this case, generalizations of Eqs. (20) and (21) for $j$th disorder realization read
$\langle 1_{c},{\bf 0}_{\rm qu}|\hat{U}(t)|1_{c},{\bf 0}_{\rm qu}\rangle_{j}=\sum_{m=1}^{N+1}\frac{\prod_{n=1}^{N}(E_{m}^{j}-\omega_{n}^{j})}{\prod_{m^{\prime}\neq m}(E_{m}^{j}-E_{m^{\prime}}^{j})}e^{-iE_{m}^{j}t},$
$\langle 0_{c},{1}_{\rm qu}^{n}|\hat{U}(t)|1_{c},{\bf 0}_{\rm qu}\rangle_{j}=\sum_{m=1}^{N+1}\frac{g_{n}\prod_{k\neq n}(E_{m}^{j}-\omega_{k}^{j})}{\prod_{m^{\prime}\neq m}(E_{m}^{j}-E_{m^{\prime}}^{j})}e^{-iE_{m}^{j}t}.$
Here $E_{m}^{j}$ are the eigenenergies which can be obtained by diagonalization of the Hamiltonian (1). The photonic population reads then
$P_{\rm ph}^{\rm av}(t)=M\big{[}|\langle 1_{c},{\bf 0}_{\rm qu}|\hat{U}(t)|1_{c},{\bf 0}_{\rm qu}\rangle_{j}|^{2}\big{]},$ (28)
while the absorption spectrum becomes
$F^{\rm av}(\omega)$ $=$ $M\Big{[}2{\Re}\Big{\{}\sum_{m=1}^{N+1}\frac{\prod_{n=1}^{N}(E_{m}^{j}-\omega_{n}^{j})}{\prod_{m^{\prime}\neq m}(E_{m}^{j}-E_{m^{\prime}}^{j})}\delta(\omega-E_{m}^{j})\Big{\}}\Big{]}$
(the superscript ”av” stays for averaged). The finite lifetimes are accounted for by replacing
$\delta(\omega-E_{m}^{j})\rightarrow\frac{1}{\pi}\frac{\Im(E_{m}^{j})}{[\omega-{\Re}(E_{m}^{j})]^{2}+\Im(E_{m}^{j})^{2}}.$
Appendix B Time evolution of the multi-D2 Ansatz
The time evolution of variational parameters $\mu_{i}$ (i.e., $A_{mn}(t)$ and $f_{mk}(t)$) is obtained by the time-dependent variational principle
$\frac{d}{dt}\frac{\partial L}{\partial\dot{\mu}_{i}^{\ast}}-\frac{\partial L}{\partial\mu_{i}^{\ast}}=0,$ (30)
where the Lagrangian $L$ reads
$L$ $=$ $\frac{i}{2}\left[\langle{\rm D}_{2}^{M}(t)|\frac{\overrightarrow{\partial}}{\partial t}|{\rm D}_{2}^{M}(t)\rangle-\langle{\rm D}_{2}^{M}(t)|\frac{\overleftarrow{\partial}}{\partial t}|{\rm D}_{2}^{M}(t)\rangle\right]$ (31)
$-\langle{\rm D}_{2}^{M}(t)|\hat{H}_{\rm HTC}|{\rm D}_{2}^{M}(t)\rangle.$
If $M$ is sufficiently large, the multi-D2 Ansatz converges to the numerically “accurate” solution to the Schrödinger equation. In terms of the multi-$\mathrm{D}_{2}$ parameters, the time evolution of the photon-mode population can be evaluated as follows
$P_{\rm ph}(t)=\sum_{m,m^{\prime}=1}^{M}A_{m1_{c}}^{*}A_{m^{\prime}1_{c}}S_{mm^{\prime}},$ (32)
where the Debye-Waller factor is given by
$S_{mm^{\prime}}=e^{\sum_{k}\left\{-\frac{1}{2}\left(\left|f_{mk}\right|^{2}+\left|f_{m^{\prime}k}\right|^{2}\right)+f_{mk}^{\ast}f_{m^{\prime}k}\right\}}.$ (33)
Using the definition of Eq. (24), we obtain the following multi-$\rm D_{2}$ expression for the absorption spectrum:
[…]
×$\times e^{\sum_{k}\left\{-\frac{1}{2}(|f_{mk}(t)|^{2}+|f_{m^{\prime}k}(0)|^{2})+f_{m^{\prime}k}^{*}(0)f_{mk}(t)\right\}}$
where $\gamma^{\prime}$ is the electronic dephasing rate.
Nonlinear spectroscopy is another important aspect in the dynamics investigation of various exciton-phonon-photon systems, as it provides valuable information on a variety of correlation functions. The information provided by the linear absorption spectra and 2D photon echo (PE) spectra gives direct knowledge on exciton-phonon interactions and on the dephasing and relaxation processes that is elusive in the output from the traditional 1D spectroscopy. In order to simulate the 2D PE spectra, we have to consider the interaction between the system and the classical light field. The corresponding Hamiltonian is given by $\hat{H}_{\rm L}=-\left(\mathbf{E}\left(\mathbf{r},t\right)\cdot\hat{\mu}_{+}+\mathbf{E}^{*}\left(\mathbf{r},t\right)\cdot\hat{\mu}_{-}\right)$, with $\mathbf{E}(r,t)$ being the time-dependent electric field of the applied pulse sequence,
$\mathbf{E}\left(\mathbf{r},t\right)$ $=$ $\mathbf{E}_{1}\left(\mathbf{r},t\right)+\mathbf{E}_{2}\left(\mathbf{r},t\right)+\mathbf{E}_{3}\left(\mathbf{r},t\right),$
$\mathbf{E}_{1}\left(\mathbf{r},t\right)$ $=$ $\mathbf{e}_{1}E_{1}\left(t-\tau_{1}\right)e^{i\mathbf{k}_{1}\cdot\mathbf{r}-i\omega_{1}t+i\phi_{1}},$
$\mathbf{E}_{2}\left(\mathbf{r},t\right)$ $=$ $\mathbf{e}_{2}E_{2}\left(t-\tau_{2}\right)e^{i\mathbf{k}_{2}\cdot\mathbf{r}-i\omega_{2}t+i\phi_{2}},$
$\mathbf{E}_{3}\left(\mathbf{r},t\right)$ $=$ $\mathbf{e}_{3}E_{3}\left(t-\tau_{3}\right)e^{i\mathbf{k}_{3}\cdot\mathbf{r}-i\omega_{3}t+i\phi_{3}},$ (35)
where $\mathbf{e}_{a}$, $k_{a}$, $\omega_{a}$, $E_{a}\left(t\right)$, and $\phi_{a}\left(a=1,2,3\right)$ denote the polarization, the wave vector, the frequency, the dimensionless envelope, and the initial phase, respectively. It is common to define the pulse arrival times in $\hat{H}_{\rm L}$ as $\tau_{1}=-T_{w}-\tau,\thinspace\tau_{2}=-T_{w},\thinspace\tau_{3}=0$ where $\tau$ (the so-called coherence time) is the delay time between the second and the first pulse, and $T_{w}$ (the so-called population time) is the delay time between third and second pulse. In the short pulse limit, we have $E_{a}\left(t\right)=E_{0}\delta\left(t\right)$. Before the optical excitation $(t\ll-T-\tau)$, the system is assumed to be in its global ground state $\left|g\right\rangle\left|0\right\rangle_{\rm ph}$. Adopting this factorized initial condition allows one to neglect correlations between the primary system and its environment. Without the contribution of excited-state absorption, the PE third polarization $P^{(3)}(t)$ can be decomposed into four nonlinear response functions $R_{1-4}$, which form the theoretical bases for the simulations of different 2D spectra. The third-order response is directly given by the response functions in the impulsive limit, which can be expressed through the multi-$\rm D_{2}$ parameters as follows sun2
$R_{1}(\tau,T_{w},t)$ $=$ $\sum_{i,j}^{M}\sum_{n,n_{1},n_{2},n_{3}}(\mathbf{e}_{4}^{*}\cdot\mathbf{\mu}_{n_{2}}^{*})(\mathbf{e}_{1}\cdot\mathbf{\mu}_{n_{3}})(\mathbf{e}_{2}^{*}\cdot\mathbf{\mu}_{n}^{*})(\mathbf{e}_{3}\cdot\mathbf{\mu}_{n_{1}})$
×$\times A_{jn_{1}n}^{*}(T_{w})A_{in_{2}n_{3}}(\tau+T_{w}+t)e^{\sum_{q}f_{jqn}^{*}(T_{w})f_{iqn_{3}}(\tau+T_{w}+t)e^{i\omega_{q}t}}$
×$\times e^{-\frac{1}{2}\sum_{q}(|f_{jqn}(T_{w})|^{2}+|f_{iqn_{3}}(\tau+T_{w}+t)|^{2})},$
$R_{2}(\tau,T_{w},t)$ $=$ $\sum_{i,j}^{M}\sum_{n,n_{1},n_{2},n_{3}}(\mathbf{e}_{4}^{*}\cdot\mathbf{\mu}_{n_{2}}^{*})(\mathbf{e}_{1}^{*}\cdot\mathbf{\mu}_{n}^{*})(\mathbf{e}_{2}\cdot\mathbf{\mu}_{n_{3}})(\mathbf{e}_{3}\cdot\mathbf{\mu}_{n_{1}})$
×$\times A_{jn_{1}n}^{*}(\tau+T_{w})A_{in_{2}n_{3}}(T_{w}+t)e^{\sum_{q}f_{jqn}^{*}(\tau+T_{w})f_{iqn_{3}}(T_{w}+t)e^{i\omega_{q}t}}$
×$\times e^{-\frac{1}{2}\sum_{q}(|f_{jqn}(\tau+T_{w})|^{2}+|f_{iqn_{3}}(T_{w}+t)|^{2})},$
$R_{3}(\tau,T_{w},t)$ $=$ $\sum_{i,j}^{M}\sum_{n,n_{1},n_{2},n_{3}}(\mathbf{e}_{4}^{*}\cdot\mathbf{\mu}_{n_{2}}^{*})(\mathbf{e}_{1}^{*}\cdot\mathbf{\mu}_{n}^{*})(\mathbf{e}_{2}\cdot\mathbf{\mu}_{n_{1}})(\mathbf{e}_{3}\cdot\mathbf{\mu}_{n_{3}})$
×$\times A_{jn_{1}n}^{*}(\tau)A_{in_{2}n_{3}}(t)e^{\sum_{q}f_{jqn}^{*}(\tau)f_{iqn_{3}}(t)e^{i\omega_{q}\left(T_{w}+t\right)}}$
×$\times e^{-\frac{1}{2}\sum_{q}(|f_{jqn}(\tau)|^{2}+|f_{iqn_{3}}(t)|^{2})},$
$R_{4}(\tau,T_{w},t)$ $=$ $\sum_{i,j}^{M}\sum_{n,n_{1},n_{2},n_{3}}(\mathbf{e}_{4}^{*}\cdot\mathbf{\mu}_{n}^{*})(\mathbf{e}_{1}^{*}\cdot\mathbf{\mu}_{n_{3}}^{*})(\mathbf{e}_{2}\cdot\mathbf{\mu}_{n_{2}})(\mathbf{e}_{3}\cdot\mathbf{\mu}_{n_{1}})$ (36)
×$\times A_{jn_{1}n}^{*}(-t)A_{in_{2}n_{3}}(\tau)e^{\sum_{q}f_{jqn}^{*}(-t)f_{iqn_{3}}(\tau)e^{-i\omega_{q}T_{w}}}$
×$\times e^{-\frac{1}{2}\sum_{q}(|f_{jqn}(-t)|^{2}+|f_{iqn_{3}}(\tau)|^{2})}.$
Here $\mathbf{e}_{4}$ refers to the polarization of the local oscillator field, and $\mathbf{\mu}_{n}$ are the transition dipole moment vectors. $A_{jn_{1}n}^{*}(t)$ represents the probability amplitude at time $t$ for the exciton at the state $\left|n_{1}\right\rangle$ with the initial state $\left|n\right\rangle$ and multiplicity $j$, and $f_{jqn}(t)$ is the corresponding phonon displacement, also starting from $\left|n\right\rangle{\exp}\left\{\sum_{q}^{N_{q}}\left[f_{jq}(0)\hat{b}_{q}^{\dagger}-f_{jq}^{*}(0)\hat{b}_{q}\right]\right\}\left|\mathbf{0}\right\rangle_{\rm ph}$.
Similarly, the higher excited-state response functions are given by sun2
$R_{1}^{*}(\tau,T_{w},t)$ $=$ $\sum_{i,j}^{M}\sum_{\mbox{\tiny$\begin{array}[]{c}nn_{1}n_{2}\\ n_{3}mm_{1}\end{array}$}}(\textbf{e}_{4}^{*}\cdot\mathbf{\mu}_{n_{1}m}^{*})(\textbf{e}_{1}^{*}\cdot\mathbf{\mu}_{n}^{*})(\textbf{e}_{2}\cdot\mathbf{\mu}_{n_{3}})(\textbf{e}_{3}\cdot\mathbf{\mu}_{n_{2}m_{1}})$
$A^{j*}_{mm(n_{1}n)}(0)A^{i}_{mm_{1}(n_{2}n_{3})}(t)e^{\sum_{q}f^{j*}_{mm(n_{1}n),q}(0)f^{i}_{mm_{1}(n_{2}n_{3}),q}(t)}$
$e^{-\frac{1}{2}\sum_{q}(|f^{j*}_{mm(n_{1}n),q}(0)|^{2}+|f^{i}_{mm_{1}(n_{2}n_{3}),q}(t)|^{2})},$
$R_{2}^{*}(\tau,T_{w},t)$ $=$ $\sum_{i,j}^{M}\sum_{\mbox{\tiny$\begin{array}[]{c}nn_{1}n_{2}\\ n_{3}mm_{1}\end{array}$}}(\textbf{e}_{4}^{*}\cdot\mathbf{\mu}_{n_{1}m}^{*})(\textbf{e}_{1}\cdot\mathbf{\mu}_{n_{3}})(\textbf{e}_{2}^{*}\cdot\mathbf{\mu}_{n}^{*})(\textbf{e}_{3}\cdot\mathbf{\mu}_{n_{2}m_{1}})$
$A^{j\prime*}_{mm(n_{1}n)}(0)A^{i\prime}_{mm_{1}(n_{2}n_{3})}(t)e^{\sum_{q}f^{j\prime*}_{mm(n_{1}n),q}(0)f^{i\prime}_{mm_{1}(n_{2}n_{3}),q}(t)}$
$e^{-\frac{1}{2}\sum_{q}(|f^{j\prime*}_{mm(n_{1}n),q}(0)|^{2}+|f^{i\prime}_{mm_{1}(n_{2}n_{3}),q}(t)|^{2})}.$
The initial amplitudes of the higher excited-state are $A^{j*}_{mm(n_{1}n)}(0)=A^{*}_{jn_{1}n}(\tau+T_{w}+t)$, $A^{i}_{mm_{1}(n_{2}n_{3})}(0)=A_{in_{2}n_{3}}(T_{w})$, $A^{j\prime*}_{mm(n_{1}n)}(0)=A^{*}_{jn_{1}n}(t+T_{w})$, and $A^{i\prime}_{mm_{1}(n_{2}n_{3})}(0)=A_{in_{2}n_{3}}(\tau+T_{w})$, and the corresponding phonon displacements are $f^{j*}_{mm(n_{1}n),q}(0)=f^{*}_{jn_{1}q}(\tau+T_{w}+t)$, $f^{i}_{mm_{1}(n_{2}n_{3}),q}(0)=f_{in_{2}q}(T_{w})$, $f^{j\prime*}_{mm(n_{1}n),q}(0)=f^{*}_{jn_{1}q}(t+T_{w})$, and $f^{i\prime}_{mm_{1}(n_{2}n_{3}),q}(0)=f_{in_{2}q}(\tau+T_{w})$.
The contributions from stimulated emission, ground-state bleach, and excited-state absorption, are expressed, respectively, as,
[…]
The correlated 2D spectrum is given by the sum of the three contributions
$S(\omega_{\tau},T_{w},\omega_{t})=$ $S_{\rm SE}(\omega_{\tau},T_{w},\omega_{t})+S_{\rm GSB}(\omega_{\tau},T_{w},\omega_{t})$ (45)
$+S_{\rm ESA}(\omega_{\tau},T_{w},\omega_{t}).$
Appendix C A pedestrian introduction to TFD
Introduced in the 1970s to study quantum mechanics at finite temperatures, the TFD method has since found applications in various problems. We combine the TFD method with the approach of multi-D2 trial states and use it to study the finite-temperature dynamics of the HTC model. In the TFD framework, to solve the Schrödinger equation under a certain initial condition, we assume that the system is in thermal equilibrium initially at a given temperature, for which the initial state can be written as
$|\psi(0)\rangle=\rho_{\mathrm{ph}}^{1/2}|I\rangle|n\rangle|\tilde{n}\rangle=|\mathbf{0}(\beta)\rangle|n\rangle|\tilde{n}\rangle.$ (46)
Here $\rho_{\mathrm{ph}}$ is the density matrix of bath DOFs, which is usually represented by a set of resonators. $\tilde{H}$ is called the tildian and represents a fictitious Hamiltonian operator that can be derived from the original operator ${H}$ using a well defined mathematical procedure. $|{\tilde{n}}\rangle$ are arbitrary basis vectors of the fictitious space. $|\mathbf{0}(\beta)\rangle$ is the temperature-dependent vacuum state and $|n\rangle(|\tilde{n}\rangle)$ is the basis vector of the excitonic part of the Hamiltonian $H(\tilde{H})$. The dimension of the TFD wave function $|\psi(t)\rangle$ in the basis set representation is $\left(N_{\mathrm{ex}}\times N_{\mathrm{ph}}\right)^{2}$ , $N_{\mathrm{ex}}\left(N_{\mathrm{ph}}\right)$ denotes the number of electron (vibrational) basis functions, and the square is due to the doubling of the DOFs. This is exactly the same dimension as the density matrix of the physical system described, and therefore the TFD can be applied directly. We note that Eq. (46) introduces a temperature effect using only the vibrational degrees of freedom, which suggests that the tilded part of the excitonic Hamiltonian $(|\tilde{n}\rangle)$ can fall. This statement was rigorously proved by Borrelli and Gelin BG21 who showed that the TFD Schrödinger equation assumes the form
$i\frac{\partial}{\partial t}|\psi(t)\rangle=\bar{H}|\psi(t)\rangle.$ (47)
The corresponding initial conditions can be written as
$|\psi(0)\rangle=\rho_{\mathrm{ph}}^{1/2}|I\rangle|n\rangle=|\mathbf{0}(\beta)\rangle|n\rangle.$ (48)
The resulting expectation $\langle Q(t)\rangle$ is consistent with the expectation calculated by the Schrödinger equation with the Hamiltonian measure $\hat{H}$ after rewriting by TFD. The redefined Hamiltonian operator $\bar{H}$ is
$\bar{H}=H-\tilde{H}_{\mathrm{ph}}.$ (49)
where ${H}$ is defined in Eq. (II), and $\tilde{H}_{\mathrm{ph}}$ is any operator acting in the tilded vibrational space. It is worth noting that the particular form of $\tilde{H}_{\mathrm{ph}}$ does not affect the expectation value $\langle Q(t)\rangle$. For simplicity, we have chosen
$\tilde{H}_{\mathrm{ph}}=\sum_{q}\omega_{q}\tilde{b}_{q}^{\dagger}\tilde{b}_{q}.$ (50)
We obtain finally the new Schrödinger equation by the Bogoliubov variation of Eq. (47)
$i\frac{\partial}{\partial t}|\varphi(t)\rangle=\bar{H}_{\theta}|\varphi(t)\rangle,\quad|\varphi(0)\rangle=|\mathbf{0}\rangle|n\rangle,$ (51)
The transformed Hamiltonian ${H}_{\theta}$ takes the form
$\bar{H}_{\theta}=$ $e^{iG}\bar{H}e^{-iG}$ (52)
$=$ $\omega_{c}{a}^{\dagger}{a}+\sum_{n=1}^{N}\left[\omega_{n}{\sigma}_{n}^{+}{\sigma}_{n}^{-}+g_{n}\left({a}^{\dagger}{\sigma}_{n}^{-}+{a}{\sigma}_{n}^{+}\right)\right]$
$+\sum_{k}\omega_{k}\left(b_{k}^{\dagger}b_{k}-\tilde{b}_{k}^{\dagger}\tilde{b}_{k}\right)$
$-\frac{\lambda}{\sqrt{N}}\sum_{k,n}\cosh\theta_{k}\omega_{k}{\sigma}_{n}^{+}{\sigma}_{n}^{-}\left(e^{ikn}b_{k}+e^{-ikn}b_{k}^{+}\right)$
$-\frac{\lambda}{\sqrt{N}}\sum_{k,n}\sin\theta_{k}w_{k}\hat{\sigma}_{n}^{+}\hat{\sigma}_{n}^{-}\left(e^{ikn}\tilde{b}_{k}^{\dagger}+e^{-ikn}\tilde{b}_{k}\right).$
Here $G$ is the Bogoliubov unitary transformation operator defined by
$G=G^{\dagger}=-i\sum_{k}\theta_{k}\left(b_{k}\tilde{b}_{k}-b_{k}^{\dagger}\tilde{b}^{\dagger}_{k}\right).$ (53)
with
$\theta_{k}=\operatorname{arctanh}\left(e^{-\beta\omega_{k}/2}\right).$ (54)
The temperature effect is thus taken into account by coupling the physical system to the fictitious tilde system through the temperature-dependent mixing angle $\theta_{k}$.
The following relations have been used:
$e^{iG}b_{k}e^{-iG}=b_{k}\cosh\left(\theta_{k}\right)+\tilde{b}_{k}^{\dagger}\sinh\left(\theta_{k}\right)$ (55)
$e^{iG}\tilde{b}_{k}e^{-iG}=\tilde{b}_{k}\cosh\left(\theta_{k}\right)+b_{k}^{\dagger}\sinh\left(\theta_{k}\right)$
$e^{iG}\left(b_{k}^{\dagger}b_{k}-\tilde{b}_{k}^{\dagger}\tilde{b}_{k}\right)e^{-iG}=b_{k}^{\dagger}b_{k}-\tilde{b}_{k}^{\dagger}\tilde{b}_{k}.$
In numerical calculations, the high-frequency modes can be dropped from the tilded Hamiltonian quantities even at room temperature owing to $\theta_{k}\ll 1,\sinh\left(\theta_{k}\right)\approx 0$, and $\cosh\left(\theta_{k}\right)\approx 1$. Therefore, there is an additional reduction in dimensionality compared to the standard Liouville’s equation.
After the introduction of the tilded phonons, the multi-D2 Ansatz with time-varying components, defined originally in Eq. (8), can be recast as
$|\tilde{\mathrm{D}}_{2}^{M}(t)\rangle=$ $\sum_{n}|n\rangle\sum_{m=1}^{M}A_{mn}(t)e^{\left(\sum_{k}f_{mk}(t){b}_{k}^{\dagger}\right)}|0\rangle_{\mathrm{ph}}$ (56)
×$\times e^{\left(\sum_{k}\tilde{f}_{mk}(t)\tilde{b}_{k}^{\dagger}\right)}|\tilde{0}\rangle_{\mathrm{ph}}.$
Here the non-normalized Bargmann coherent state is adopted. The advantage is that it does not need the additional conjugate coherent state parameter space in the calculations, which doubles the computing efficiency in comparison to the case of the normalized coherent state. $|\tilde{\rm{D}}_{2}^{M}(t)\rangle$ in Eq. (56) can be used to solve Eq. (51).
According to the variational parameters obtained by the multi-$\mathrm{D}_{2}$ method, the time evolution of the population of the photon modes can be expressed in the following form:
$P_{\mathrm{ph}}(t)=\sum_{m,m^{\prime}=1}^{M}A_{m1_{c}}^{*}A_{m^{\prime}1_{c}}S_{mm^{\prime}}.$ (57)
where the Debye-Waller factor is
$S_{mn^{\prime}}=\exp\left[\sum_{k}f_{mk}^{*}(t)f_{m^{\prime}k}(t)+\sum_{k}\tilde{f}_{mk}^{*}(t)\tilde{f}_{m^{\prime}k}(t)\right].$ (58)
The multi-$\mathrm{D}_{2}$ expression for the absorption spectrum is
[…]
×$\times e^{\left[{\sum_{k}f_{m^{\prime}k}^{*}(0)f_{m^{\prime}k}(t)+\sum_{k}\tilde{f}_{m^{\prime}k}^{*}(0)\tilde{f}_{m^{\prime}k}(t)}\right]}.$